### v2

# bayesNMF: Fast Bayesian Poisson NMF with Automatically Learned Rank Applied to Mutational Signatures

Abstract
Bayesian Non-Negative Matrix Factorization (NMF) is a method of interest in fields including genomics, neuroscience, and audio and image processing. Bayesian Poisson NMF is of particular importance for counts data, for example in cancer mutational signature analysis. However, MCMC methods for Bayesian Poisson NMF require a computationally intensive augmentation. Further, identifying latent rank is necessary, but commonly used heuristic approaches are slow and potentially subjective, while methods that learn rank automatically are unable to provide posterior uncertainties. In this paper, we introduce bayesNMF, a computationally efficient Gibbs sampler for Bayesian Poisson NMF. Metropolis-Hastings steps are used to avoid augmentation, where full conditionals from a Normal-likelihood NMF is used as geometry-informed, high-overlap proposals. We additionally define sparse Bayesian factor inclusion (SBFI) as a method to identify rank automatically while providing posterior uncertainty quantification. We provide an open-source R software package with all of the models and plotting capabilities demonstrated in this paper on GitHub at jennalandy/bayesNMF, and supplemental materials are available online. Although our applications focus on cancer mutational signatures, our software and results can be extended to any use of Bayesian Poisson NMF.
Keywords Non-negative matrix factorization $\cdot$ Efficient Bayesian computation $\cdot$ Gibbs sampling $\cdot$ Mutational signatures analysis $\cdot$ Poisson augmentation
1 Introduction
Non-Negative Matrix Factorization (NMF) is a widely used unsupervised technique for dimension reduction and intuitive parts-based representation [Lee and Seung, 1999]. In cancer genomics, Poisson NMF can be applied to mutation counts to model mutational processes with latent factors, assuming that multiple processes contribute additively to tumor genomes [Alexandrov et al., 2013]. Bayesian Poisson NMF allows for uncertainty quantification and the integration of prior knowledge [Cemgil, 2009, Schmidt et al., 2009].
The computational cost of MCMC methods (e.g., Gibbs samplers) for Bayesian NMF grows with the number of samples $G$ (columns), latent factors $N$, and variables $K$ (rows). With a Normal likelihood, conditional updates take the form of standard distributions, requiring $(K+G)\times N$ Gibbs updates per iteration, plus updates for residual variance [Schmidt et al., 2009]. In contrast, Bayesian NMF with a Poisson likelihood involves a data augmentation step, called Poisson augmentation, that decomposes each observed count into latent factor-specific counts, introducing $K\times G\times N$ auxiliary latent variables. This results in $(K+G+KG)\times N$ Gibbs updates [Cemgil, 2009], leading to a substantially higher computational cost. With growing sources of publicly available data for the analysis of mutational signatures [Consortium, 2010, Weinstein et al., 2013, Consortium, 2020] and a growing reference set of previously discovered signatures [Alexandrov et al., 2020], computational efficiency is a key consideration for work in this field.
In addition, identifying latent rank is a fundamental aspect of any factor model. Most Bayesian NMF methods take a heuristic approach: they fit a separate model for each rank and optimize regularized metrics such as the Bayesian information criterion (BIC) [Rosales et al., 2016, Gori and Baez-Ortega, 2018] or determine the “elbow” of metrics such as mean squared error (MSE) or cosine similarity [Gori and Baez-Ortega, 2018, Islam et al., 2022]. This is a computationally intensive process and requires post hoc, and potentially subjective, model selection. Existing methods that automatically learn rank as part of the Bayesian model with automatic relevance determination (ARD) do so by optimizing the posterior numerically, which means they are fast but result in point estimates without posterior uncertainties [Kasar et al., 2015]. Another recent method, compressive NMF [Zito and Miller, 2024, CompNMF], is able to learn rank automatically while maintaining posterior uncertainty, but it does not ease the computational burden of Poisson augmentation.
Table 1 summarizes the relative strengths of existing methods for Poisson Bayesian NMF and identifies the gap that this work addresses: we introduce a Gibbs sampler for Bayesian Poisson NMF which 1) is computationally efficient by avoiding Poisson augmentation, 2) automatically determines latent rank as a part of the Bayesian model, and 3) reports posterior uncertainties. We avoid the computationally intensive Poisson augmentation through Metropolis Hastings sampling with high-overlap, geometry-informed proposals derived from a Normal-likelihood Bayesian NMF. We show the accuracy and efficiency of this approach through simulation studies and apply it to 32 cancer types from the Pan-Cancer Analysis of Whole Genomes (PCAWG) database [Consortium, 2020]. We also provide an open-source R software package, bayesNMF, which is available on GitHub at jennalandy/bayesNMF.
2 Background
Software Automatically Posterior Avoids Poisson
Learns Rank Uncertainty Augmentation
bayesNMF (this work) $\checkmark$ $\checkmark$ $\checkmark$
SignatureAnalyzer (2015) $\checkmark$ $\checkmark$
SigneR (2016) $\checkmark$
SigFit (2018) $\checkmark$
CompNMF (2024) $\checkmark$ $\checkmark$
Table 1: Software for Bayesian NMF applied to de-novo mutational signature analysis: SignatureAnalyzer [Kasar et al., 2015, Kim et al., 2016], SigneR [Rosales et al., 2016], SigFit [Gori and Baez-Ortega, 2018], and CompNMF [Zito and Miller, 2024]. An ideal software would learn rank automatically, report posterior uncertainty, and avoid computationally intensive Poisson augmentation.
2.1 Non-Negative Matrix Factorization (NMF) and Bayesian NMF
NMF is an unsupervised method to decompose non-negative data $M$ ($K$ variables by $G$ samples) into two lower-rank non-negative matrices, $P$ and $E$ [Paatero and Tapper, 1994, Lee and Seung, 1999]. NMF represents each observation of the data matrix, or column of $M$, as a linear combination of latent factors, or columns of $P$, with weights in $E$:
$M_{kg}$ $=\sum_{n=1}^{N}P_{kn}E_{ng},\hskip 11.38109pt\mbox{i.e}\;\;M=PE,\hskip 11.38109ptM\in\mathbb{R}_{\geq 0}^{K\times G},P\in\mathbb{R}_{\geq 0}^{K\times N},E\in\mathbb{R}_{\geq 0}^{N\times G}.$
NMF is typically fit by minimizing a reconstruction error, often the generalized Kullback-Leibler (KL) divergence between $M$ and $PE$ (Equation 1) or the Frobenius norm of MPEM-PE (Equation 2). For each, Lee and Seung [2000] adapt gradient descent by deriving a step size that preserves nonnegativity, resulting in multiplicative update algorithms.
$KL(M||PE)$ $=\sum_{kg}\left(M_{kg}log\frac{M_{kg}}{(PE)_{kg}}-M_{kg}+(PE)_{kg}\right)$ (1)
[…]
Minimizing KL-divergence is equivalent to maximizing a Poisson likelihood on $M$ centered at $PE$ [Cemgil, 2009]. Similarly, minimizing the Frobenius norm is equivalent to maximizing a Normal likelihood on $M$ centered at $PE$ [Schmidt et al., 2009].
Bayesian approaches add priors on $P$ and $E$ [Cemgil, 2009, Schmidt et al., 2009]. To sample Bayesian Poisson NMF posterior distributions with standard MCMC methods, Poisson augmentation is necessary to achieve easily-sampled full conditional distributions of $P$ and $E$ [Cemgil, 2009]. This introduces latent variables $Z_{kng}$, defined as the part of the $M_{kg}$ count attributed to latent factor $n$. Each $Z_{kng}$ follows a Poisson distribution centered at $P_{kn}E_{ng}$ and satisfies $M_{kg}=\sum_{n}Z_{kng}$. This imposes $K\cdot N\cdot G$ additional Gibbs updates at each iteration, increasing the computational requirements of Bayesian Poisson NMF far beyond that of Bayesian Normal NMF.
2.2 Cancer Mutational Signatures
Cancer is driven by genetic changes, including somatic DNA mutations, which can be caused by multiple processes. Understanding which mutational process activities are reflected in a tumor genome can improve our understanding of cancer development and answer questions about subtyping, prognosis, and treatment [Nik-Zainal et al., 2012a, b, Alexandrov et al., 2013, Rosales et al., 2016]. A computationally-derived mutational signature models the specific patterns produced by a mutational process. Examples include DNA damage repair deficiencies [Zámborszky et al., 2017], deamination of 5-methylcytosine [Nik-Zainal et al., 2012a], tobacco smoke [Alexandrov et al., 2013, Nik-Zainal et al., 2015], and ultraviolet radiation [Nik-Zainal et al., 2015, Saini et al., 2016, Hayward et al., 2017]. The Catalog of Somatic Mutations in Cancer (COSMIC) [Tate et al., 2019] provides a high-confidence set of reference signatures. These are an important reference, but cannot be treated as ground truth, as they were estimated from data.
[…]
Mutational signature analysis is a rapidly growing field driven by an increasing number of publicly available databases, including COSMIC [Tate et al., 2019], mSignatureDB [Huang et al., 2018], the Pan-Cancer Analysis of Whole Genomes (PCAWG) [Consortium, 2020], and the compendium of Mutational Signatures of Environmental Agents [Kucab et al., 2019]. With increased data sizes and availability, computational efficiency must be carefully considered when making modeling and implementation choices.
3 Methodology
3.1 MH-within-Gibbs for Computational Efficiency
We introduce two efficient models for Poisson Bayesian NMF. The first uses a Truncated Normal (TN) prior on each element of $P$ and $E$:
Likelihood: $\quad M_{kg}\sim\text{Poisson}((PE)_{kg})$ (3)
Priors: $\quad P_{kn}\sim\text{TN}(\mu^{P}_{kn},\sigma^{P}_{kn},0,\infty),\quad E_{ng}\sim\text{TN}(\mu^{E}_{ng},\sigma^{E}_{ng},0,\infty).$ (4)
Specifications of hyperprior distributions and hyperparameters can be found in Appendix A.
The target distributions $f$ for Gibbs updates of $P$ and $E$ are not easily sampled. Using $P_{kn}$ as an example, let $P^{*}_{kn}$ be a potential new value, and let $P^{*}$ be the $P$ matrix with element $k,n$ replaced by $P^{*}_{kn}$. Each term $(P^{*}E)_{kg}$ depends on $P^{*}_{kn}$ for all $n$, linking likelihoods across $n$, and preventing factorization into a standard family:
$f(P_{kn}^{*}|P_{-kn},E,M)$ $\propto p_{\text{TN}}(P_{kn}^{*})\prod_{g}p_{\text{Poisson}}\left(M_{kg}|\lambda=(P^{*}E)_{kg}\right)$ (5)
$\propto p_{\text{TN}}(P^{*}_{kn})e^{-\sum_{g}(P^{*}E)_{kg}}\prod_{g}(P^{*}E)_{kg}^{M_{kg}}.$
Poisson augmentation resolves this in previous methods [Cemgil, 2009], but at high computational cost. Instead, we use Metropolis-Hastings (MH) steps for Gibbs updates of the $P$ and $E$ matrices. We refer to models using this approach as “Poisson+MH” from here on. We construct proposal distributions $g$ from the full conditional form of a Normal Bayesian NMF model with identical Truncated Normal priors to our target model:
$g(P^{*}_{kn}$ $|P_{-kn},E,M)\propto p_{\text{TN}}(P^{*}_{kn})\prod_{g}p_{\text{Normal}}\left(M_{kg}|\mu=(P^{*}E)_{kg},\sigma^{2}=(PE)_{kg}\right)$ (6)
$\sim\text{Truncated Normal}\left(P^{*}_{kn}|\mu=m(M,P,E,\theta),\sigma=s(M,P,E,\theta),0,\infty\right)$
where $\theta$ holds prior and hyperprior parameters. The Normal model is never sampled directly. Specifications of the location $m(\cdot)$ and scale $s(\cdot)$ functions are in Appendix A.
Instead of modeling variances $\sigma^{2}_{kg}$ as would be required in a Normal Bayesian NMF, we substitute it with $(PE)_{kg}$ in the proposal. This allows the proposal to reflect the Poisson mean-variance relationship of the target model without additional parameters. This only affects the proposal $g$, which needs to be a proper distribution but does not need to be an actual full conditional to maintain detailed balance.
After drawing a proposed value $P^{*}_{kn}$, the acceptance ratio can be computed as a ratio of likelihoods because priors cancel out from both the target and proposal distributions:
$a^{P}_{kn}$ $=\frac{f(P^{*}_{kn}|P_{-kn},E,M,\theta)g(P_{kn}|P^{*},E,M,\theta)}{f(P_{kn}|P_{-kn},E,M,\theta)g(P^{*}_{kn}|P,E,M,\theta)}$ (7)
$=\prod_{g}\frac{p_{\text{Poisson}}\left(M_{kg}|\lambda=(P^{*}E)_{kg}\right)\cdot p_{\text{Normal}}\left(M_{kg}|\mu=(PE)_{kg},\sigma^{2}=(P^{*}E)_{kg}\right)}{p_{\text{Poisson}}\left(M_{kg}|\lambda=(PE)_{kg}\right)\cdot p_{\text{Normal}}\left(M_{kg}|\mu=(P^{*}E)_{kg},\sigma^{2}=(PE)_{kg}\right)}.$ (8)
For a fixed rank, the Normal and Poisson Bayesian NMF models have identical MAP solutions (up to scaling and permutation symmetries of NMF [Lee and Seung, 1999, Donoho and Stodden, 2003]). Consequently, the Normal-based proposal captures the local curvature of the Poisson posterior near its mode, ensuring high overlap. This proposal is analogous to gradient-based methods (e.g., Metropolis-adjusted Langevin [Roberts and Tweedie, 1996] or Hamiltonian Monte Carlo [Neal et al., 2011]), but we do not compute gradients explicitly.
The second model we introduce uses an Exponential prior. The proposal still takes the form of a Truncated Normal distribution, though functions $m(\cdot)$ and $s(\cdot)$ change. In general, the Poisson+MH approach can only be used when the prior (which we assume is shared by the target and the proposal) leads to Normal Bayesian NMF full conditionals that are of known families, meaning Gamma priors cannot be used with this setup. The limitation is practical rather than fundamental—other proposal families could in principle accommodate Gamma priors but are not considered here.
3.2 Sparse Bayesian Factor Inclusion (SBFI) for Learning Rank
Previous Bayesian work by Grabski et al. [2023a, b] in the multi-study setting learns factor sharing patterns and total latent dimension using a binary factor inclusion matrix $A$. We adapt this approach to learn rank automatically in our single-study setting and introduce a new prior and hyperprior structure such that the induced prior on the expected rank is approximately uniform. We term this approach Bayesian factor inclusion (BFI).
BFI uses a diagonal binary factor inclusion matrix, $A$, where $A_{nn}=1$ if latent factor $n$ is included in the model, and 0 if excluded, redefining NMF as
$\mathbb{E}[M_{kg}]=\sum_{n=1}^{N}P_{kn}A_{nn}E_{ng},\quad\quad\mbox{i.e.}\quad\quad\mathbb{E}[M]=PAE$
with latent rank $N^{\prime}=\sum_{n=1}^{N}A_{nn}$. When $A_{nn}=0$, hyperpriors of $P$ and $E$ retain some information on the excluded signature, reducing the extent of label switching.
Each factor inclusion indicator $A_{nn}$ has a Bernoulli prior. The prior probability of factor inclusion $q$ is determined by a random variable $R$, which represents the expected rank and has a uniform hyperprior between 0 and $N$. The value $q$ is bounded away from 0 and 1 to prevent the Gibbs sampler from getting stuck at the endpoint ranks 0 or $N$:
$A_{nn}\sim\text{Bernoulli}(q),\quad\quad q=R/N\text{, truncated to }q\in\{0.4/N,1-0.4/N\}$
$p(R=r)=1/(N+1)\text{ for }r=0,...,N.$
In practice, we set $N$ to a large number. By learning the posterior distribution of $A$, we also learn the implied posterior distribution of the latent rank $N^{\prime}=\sum_{n=1}^{N}A_{nn}$.
Including $A$ in the model induces a highly multimodal posterior. We adapt the tempering approach used in Grabski et al. [2023b] for the Gibbs updates of $A$ and $R$, where the likelihood term is raised to a temperature power $\gamma$: $A_{nn}\sim p(A_{nn})p(M|A_{nn},...)^{\gamma}$ and $R\sim p(R)p(M|R,...)^{\gamma}$. The temperature $\gamma$ progresses from 0 to 1 over the course of a burn-in period. Early samples are from the prior alone, allowing better exploration of the parameter space, and once $\gamma=1$, samples are from the full conditional. No tempered samples are used for posterior inference.
We also introduce Sparse Bayesian factor inclusion (SBFI), which replaces the likelihood in the Gibbs update of $A_{nn}$ with $e^{-\frac{1}{2}\text{BIC}}$ to encourage sparsity in the latent rank:
BIC $=-2\cdot\text{log}\left[p_{\text{Poisson}}(M|A_{nn}=a,...)\right]+\text{log}(G)\cdot(K+G)N^{(a)}$
[…]
where $N^{(a)}=\sum_{n^{\prime}}A_{n^{\prime}n^{\prime}}$ when $A_{nn}=a$, $G$ is the sample size, and $(K+G)N^{(a)}$ is the number of functional parameters in $P$ and $E$. Expanding $e^{-\frac{1}{2}\text{BIC}}$ shows that, once $\gamma=1$, SBFI is equivalent to a Gibbs update under the following sparse prior on $A_{nn}$:
$p_{\text{SBFI}}(A_{nn}=1)$ $=\frac{qG^{-\frac{1}{2}(K+G)}}{qG^{-\frac{1}{2}(K+G)}+(1-q)}.$
3.3 Baseline Models
For comparison, we implement standard Gibbs samplers for (i) Poisson Bayesian NMF with Exponential and Gamma priors (using Poisson augmentation), and (ii) Normal Bayesian NMF with Exponential and Truncated Normal priors. Standard Poisson models serve as direct baselines, against which our novel Poisson+MH models reach equivalent solutions at reduced computational cost. The Normal models provide motivation and validation of our Normal model-based proposals within the Poisson+MH framework, as they converge to the same modes as the standard Poisson models. All models are summarized in Table 2.
The Poisson-Gamma model is available in SigneR [Rosales et al., 2016] while Truncated Normal and Exponential priors as well as Normal likelihoods are available in SignatureAnalyzer [Kasar et al., 2015, Kim et al., 2016]. We implement our own samplers to enable direct models comparisons where all settings but likelihoods and priors are held constant. We do not implement a standard Gibbs sampler for the Poisson-Truncated Normal model because the full conditionals are not from distributions that can be directly sampled, even with Poisson augmentation. The Poisson-Gamma model is the closest comparison.
To evaluate the ability of SBFI to learn rank, we compare it to minBIC, a procedure which fits a separate Poisson+MH sampler for each candidate rank and selects the one minimizing BIC. We also compare it to automatic relevance determination as implemented in SignatureAnalyzer, which learns rank but does not report posterior uncertainty.
Likelihood Prior Sampler Purpose Reference
Poisson Truncated Normal MH-within-Gibbs This Work -
Poisson Exponential MH-within-Gibbs This Work -
Poisson Gamma Standard Gibbs Baseline SigneR
Poisson Exponential Standard Gibbs Baseline SignatureAnalyzer
Normal Truncated Normal Standard Gibbs Motivation SignatureAnalyzer
Normal Exponential Standard Gibbs Motivation SignatureAnalyzer
Table 2: List of models implemented in bayesNMF. MH-within-Gibbs samplers are the new contribution of this work. Standard Poisson samplers serve as a baseline for comparison to validate MH-within-Gibbs models. Standard Normal samplers serve as motivation for the proposal distributions of MH-within-Gibbs models. References refer to other software implementing the same likelihood-prior combinations, though they are re-implemented in bayesNMF to allow for direct comparisons. (SignatureAnalyzer [Kasar et al., 2015, Kim et al., 2016], and SigneR [Rosales et al., 2016].)
3.4 Inference
Inference is performed on the final 1000 posterior samples. When learning rank, we first compute $\hat{A}$, the maximum a-posteriori (MAP) estimate of the signature inclusion matrix, at its mode. We then subset inference samples to those that match this MAP.
To summarize the inference samples, we next remove the scale indeterminacy of NMF by normalizing each iteration’s $P^{(i)}$ and $E^{(i)}$ matrices for columns of $P^{(i)}$ to sum to one. Then, final estimates $\hat{P}$ and $\hat{E}$ and 95% credible intervals are element-wise means and 2.5th and 97.5th percentiles, respectively. The estimated reconstruction $\hat{M}$ is the product $\hat{P}\hat{A}\hat{E}$.
To compare results with a reference $P$, we assign each inference samples’s signature matrix $P^{(i)}$ to $P$ with the Hungarian algorithm [Kuhn, 1955]. This assumes each reference signature may only be assigned to a single estimated signature, and vice versa. We ensemble assignments over inference samples with majority voting (details in Appendix B.4).
4 Implementation and R Software Package
4.1 Implementation
The bayesNMF R software package implements an automated convergence detection algorithm and reports convergence metrics in a log file and with trace plots (Figure 1A). This makes the software simple to use, reduces subjectivity of sampler stopping, and ensures all models reach a similar degree of convergence so results can be compared fairly. Full details of model specifications, Gibbs updates, and hyperparameters are available in Appendix A. Details of convergence, inference, and reference assignment are in Appendix B.
4.2 Modeling Capabilities
The bayesNMF R software package allows users to fit all models described in this paper with the bayesNMF function. Model specifications can be adjusted by the likelihood, prior, and MH parameters. The rank can be a fixed value or a range vector, in which case rank_method specifies whether minBIC, SBFI, or BFI is used to learn rank. Users are also able to set hyperprior parameters or specify initial values of parameters or prior parameters.
4.3 Reference Comparison and Visualization Capabilities
Figure 1: Illustration of software package capabilities using bayesNMF Poisson-Truncated Normal+MH SBFI on simulated data. A. Posterior diagnostic traceplots. B. Reference assignment using posterior ensemble with majority voting. C. Visualization suite, including similarity heatmaps, contribution summaries, and reconstructed signatures (bar chart of aligned reference, points for final estimates, and error bars for 95% credible intervals).
The bayesNMF R package allows users to visualize results and optionally compare them to a set of reference signatures (the default are COSMIC v3.3.1 SBS signatures [Tate et al., 2019]). Figure 1B provides an example output of the reference assignment.
The plot function creates multiple visualizations comparing the sampler results to the reference. Figure 1C shows a cosine similarity heatmap between estimated signatures and assigned references, a dot plot highlighting the median number of mutations attributed to each signature and posterior mean cosine similarity, as well as estimated mutational signatures with posterior uncertainty. Additional functions, plot variations, and a visual diagnostic for label switching are documented in the package vignettes and in Appendix B.5.
Figure 2: Relative performance of Bayesian NMF models. Grey fill represents Gamma prior on standard Poisson models, but Truncated Normal prior on Poisson+MH or Normal models. White fill always indicates Exponential priors. A. Agreement of $\hat{E}$ (left) and $\hat{P}$ (right) between models as signature-wise minimum cosine similarities. Top: standard Poisson vs Normal; Middle: standard Poisson vs Poisson+MH; Bottom: standard Poisson vs standard Poisson (null case). Excluding 0 to 7 values per density for similarity <<0.95 for visual clarity (see Appendix Table C.4). B. Efficiency gain of Poisson+MH relative to standard Poisson. C. Performance of novel Poisson+MH models.
5 Simulation Studies
5.1 Metrics
Before computing metrics, we align signatures as described in Section 3.4. SignatureAnalyzer does not report posterior samples so results are aligned on final estimates only.
When rank is known and two models are compared to one another, we report the cosine similarity of the worst-aligned signature, as well as that of the worst-aligned exposure profile. If the two solutions are identical, both values will be exactly 1. We also report relative efficiency as the ratio of total runtime between two models.
When rank is known and a model is compared to ground truth, we first evaluate NMF solutions in terms of their ability to reconstruct the original data matrix using RMSE and KL-Divergence between $M$ and $\hat{M}$. We then measure their ability to reconstruct ground truth latent factors using the cosine similarity of the worst-reconstructed signature. Finally, we report total runtime in minutes to assess computational efficiency.
When learning rank, we evaluate NMF solutions by the bias of estimated rank as well as post-alignment sensitivity and precision. Sensitivity is the proportion of true signatures that have an estimated match of cosine similarity >>0.9. It is penalized by underestimating rank because at least one true signature will be without an estimated match. Precision is the proportion of estimated signatures that have a true match. It is penalized by overestimating rank. It may also be penalized by underestimating rank if estimated signatures are combinations of true signatures, but not if one or more true signatures are simply excluded.
5.2 Poisson+MH efficiently samples the correct posterior
In this section, we further justify our Normal model-based proposals, and we show that Poisson+MH models more efficiently sample the same posterior as standard Poisson models.
We consider true latent ranks $N\in[2,4,8,16]$ and sample sizes $G\in[16,32,64,128]$ such that $G\geq 8N$. For each combination, 10 datasets are generated. $N$ signatures are sampled from COSMIC SBS v3.3.1 [Tate et al., 2019], each normalized to sum to 1. The expected number of mutations for a sample $g$ is $m_{g}\sim\text{Negative Binomial}(r\cdot N,p)$}, with $r=111.11$ and $p=0.1$. This yields $E[m_{g}]=1000\cdot N$, or $1000$ mutations per signature per sample, on average, so that each simulation provides comparable relative power to learn each signature. Columns of the exposures matrix, $E$, are generated $E_{g}\sim\text{Multinomial}(m_{g},\mathbf{p})$, where $\mathbf{p}\sim\text{Dirichlet}(\mathbf{1}_{N})$. Mutational counts are simulated as $M_{kg}\sim\text{Poisson}(\lambda_{kg}=(PE)_{kg})$.
Figure 2A compares point estimates. In all comparisons, agreement is >>0.99 in at least 90% of cases for $\hat{E}$ and 84% for $\hat{P}$. The top row shows Normal Bayesian NMF results in point estimates nearly identical to Poisson Bayesian NMF, justifying our Normal model-based MH proposals. The middle row shows the addition of MH steps does not alter point estimates of Poisson Bayesian NMF. The bottom row is a null case, comparing two standard Poisson samplers, to show that some variation is expected due to the randomness of Gibbs samplers. Along with aligned point estimates, the addition of MH steps increases effective sample sizes (Appendix C.1) and does not substantially change credible interval widths (Appendix C.3). Together, these results show the correctness of our novel Poisson+MH samplers in that they explore the same posterior space as the standard Poisson samplers.
Figure 2B reports relative efficiency of Poisson+MH compared to standard Poisson samplers. Poisson+MH models are 3-30x faster, and the gap grows with dimensionality.
Figure 2C reports metrics of the Poisson+MH samplers as a baseline for the comparisons of Figures 2A-B. Metrics for all models are available in Appendix Figure C.1.
5.3 SBFI learns rank
In this section, we demonstrate the utility of SBFI for automatically learning rank. SBFI performs comparably to, or better than, SignatureAnalyzer’s automatic relevance determination (ARD) and reports posterior uncertainty. Our recommendation is bayesNMF with SBFI for Truncated Normal priors or bayesNMF with minBIC for Exponential priors.
Figure 3: Rank bias, precision, sensitivity, and time (log scale) of rank learning approaches: bayesNMF with SBFI (upwards triangle) and minBIC (square), as well as SignatureAnalyzer with ARD (downwards triangle), each with Truncated Normal (filled) and Exponential (empty) priors. Plotted are large shapes for medians on top of lighter boxplots and jittered outliers. Full results for all ranks between 1 and 20 are available in Appendix Figure C.3. Sensitivity is the proportion of true signatures for which there is an estimated signature with cosine similarity >>0.9. Precision is the proportion of estimated signatures for which there is a true signature with cosine similarity >>0.9.
Data are simulated as in Section 5.2, but with $G=64$ and true ranks $N$ between $1$ and $20$. 10 datasets are simulated per $N$ such that no two included signatures have cosine similarity $\geq$0.8. We note that lower precision and sensitivity at the larger ranks may reflect the difficulty of the problem rather than flaws of the methods considered.
Figure 3 reports rank bias, precision, sensitivity, and runtime for each model as true rank increases along the x-axis. For a Truncated Normal / L2 prior (filled shapes), all approaches estimate rank very well, but with different error patterns. SignatureAnalyzer overestimates low and high ranks, while SBFI and minBIC both underestimate high ranks. At high ranks, SBFI and minBIC perform best in terms of precision and sensitivity, meaning that the signatures our models do estimate are correct, even though some are missed, while SignatureAnalyzer yields some noisy or split signatures even if it learns rank correctly. For an Exponential / L1 prior (empty shapes), the sparsity penalty of SBFI may be too strong, as SBFI underestimates high ranks by a wide margin (up to 8 signatures under for $N=20$). However, minBIC still outperforms SignatureAnalyzer in terms of precision and sensitivity.
With our Poisson+MH samplers, we are able to report posterior uncertainty at the cost of 12 additional minutes over SignatureAnalyzer (4 vs 16 minutes). This would take multiple hours using SBFI with standard Poisson samplers. Although choosing rank by optimizing BIC is not a new idea, our Poisson+MH samplers also allow the minBIC strategy to complete much faster: the minBIC approach takes about two hours with Poisson+MH samplers, but would take over 20 hours with standard Poisson samplers.
6 Analysis of SBS Signatures in PCAWG Data
6.1 Data
In this section, we analyze the Pan-Cancer Analysis of Whole Genomes (PCAWG) mutational counts data for whole genome sequencing (WGS) 96-alphabet single base substitution (SBS) mutation classification Consortium [2020]. This is publicly available from the International Cancer Genome Consortium’s (ICGC) Accelerating Research in Genomic Oncology (ARGO) data portal [Zhang et al., 2019, access link]. This includes 37 datasets, each for a different histology group. We excluded histology groups with 10 or fewer samples, as well as benign (non-cancerous) histology groups, leaving 32 datasets for analysis.
We exclude hypermutated samples from all analyses. The inclusion of hypermutated samples can result in signature bleeding, where mutations in non-hypermutated samples are misattributed to hypermutation-specific signatures [Maura et al., 2019, Alexandrov et al., 2020]. For each histology group, we identify hypermutated samples by assuming the total mutation count per sample follows a Negative Binomial mixture model, with at least one component for non-hypermutated samples, and possibly one or more components for hypermutated samples (details in Appendix D.1). This removes 65 samples total (<<3% of data), with at most 13 samples removed per histology group in Skin-Melanoma.
6.2 Methods
Because ground truth is unknown for real data applications, rigorous conclusions can only be made in the context of simulated data as in Section 5. However, we wish to illustrate our method and compare it to others in a real world setting. We compare our results to SignatureAnalyzer and show our approach results in similar biological interpretations.
We use Poisson-Truncated Normal+MH SBFI as this specification is the closest to the Gamma priors used in most mutational signatures literature. We fit a separate sampler to each histology group, learning ranks $\in[1,20]$. For comparison, we also fit SignatureAnalyzer with a Poisson objective and L2 priors (equivalent to Truncated Normal priors). Estimated signatures are assigned to the COSMIC reference as described in Section 3.4.
Figure 4: Results of bayesNMF+SBFI (upwards triangle) and SignatureAnalyzer+ARD (downwards triangle) on PCAWG histology groups. Skin Melanoma is excluded for visual clarity because SignatureAnalyzer estimates a high rank of 76, (bayesNMF estimates a rank of 10). A) COSMIC reference signatures aligned to $\hat{P}$. Reference signatures not found with either method are excluded. B) Latent ranks estimated by bayesNMF+SBFI and SignatureAnalyzer+ARD. Exact values in Appendix Table D.5.
6.3 Results
Figure 4A shows a high degree of overlap in the signatures discovered by bayesNMF+SBFI and SignatureAnalyzer+ARD. These results mirror many trends from Alexandrov et al. [2020], including the presence of SBS1, SBS5, and SBS40 in nearly every cancer type. More detailed results including the median number of mutations attributed to each signature in each histology group by each method can be found in Appendix Figure D.2.
A clear difference between methods is the estimated rank—bayesNMF+SBFI typically results in a sparser solution, with SignatureAnalyzer+ARD frequently estimating twice the number of signatures (Figure 4B). Based on our simulation results, because bayesNMF+SBFI with Truncated Normal priors consistently achieved higher sensitivity and precision, we speculate that results here of SignatureAnalyzer with L2 priors may include spurious components (especially for Skin Melanoma, where SignatureAnalyzer estimates a rank of 76), making the resulting signatures less biologically interpretable.
In most cases where bayesNMF recovered a signature not reported by SignatureAnalyzer (Figure 4A upwards triangles), the discrepancy can be partly explained by posterior uncertainty. These signatures have lower cosine similarity with their assigned reference signatures and a lower proportion of assignment votes, indicating uncertainty in assignment (Appendix Figure D.3). However, of the 34 signatures of this type, 28 (82%) had a SignatureAnalyzer-derived signature among the candidates receiving assignment votes, just not the majority required for posterior ensemble assignment. Therefore, the signature estimated by SignatureAnalyzer can be viewed as consistent with one plausible posterior mode inferred by bayesNMF. There are 6 signatures still unique to bayesNMF after considering this aspect, of which many match previous discoveries in the COSMIC database: SBS5 in CNS pilocytic astrocytoma and pancreatic adenocarcinoma, SBS1 in lymphocytic leukemia (CLL), and SBS95 in liver hepatocellular carcinoma (HCC). Some are not in the COSMIC database, but have been previously reported in the literature: SBS36 in prostate adenocarcinoma [Jiang et al., 2020]. However, there is no reference for SBS9 in esophageal adenocarcinoma, although it has been found in esophageal small cell carcinoma.
7 Discussion
In this paper, we introduced bayesNMF, a novel computationally efficient Gibbs sampler for Bayesian Poisson NMF and an automatic rank-learning approach. Through simulation studies and a large-scale data analysis, we demonstrated that bayesNMF achieves comparable accuracy and uncertainty quantification to standard Bayesian Poisson NMF at a fraction of the computational cost. All methods are available in an open-source R software package on GitHub at jennalandy/bayesNMF.
The Metropolis-Hastings steps are the key to bayesNMF’s computational efficiency. In future work, these updates could easily be used in conjunction with other approaches, including compressive NMF [Zito and Miller, 2024] or multistudy NMF [Grabski et al., 2023b]. This concept of model-paired proposals may also be generalizable to other hierarchical count models that rely on Poisson augmentation, for example Poisson regression with random effects.
We recognize a potential limitation of our assumption that each estimated signature must be assigned to a unique reference signature. In some cases, precision reported in Figure 3 could be higher if multiple estimated signatures were allowed to map to the same reference. Similarly, some SignatureAnalyzer-estimated signatures in the PCAWG analysis that show cosine similarity to their assigned references (Appendix Figure D.2) may have achieved higher similarity if duplicate assignments were permitted. This is only a limitation of our downstream analysis and visualization tools, not of the models themselves.
In summary, bayesNMF offers a powerful toolbox for exploring complex mutational landscapes, especially with its ability to report posterior uncertainty while learning rank as part of the Bayesian model. The open source R software package encourages usability of these methods and reproducibility of our results. As cancer genomics databases expand, bayesNMF’s efficiency and flexibility make it well suited to drive further research.
8 Acknowledgments
The authors gratefully acknowledge that this work was supported by the NIH-NIGMS training grant T32GM135117, NIH-NCI grant 5R01 CA262710-03, and NSF-DMS grant 2113707.
Supplementary Material
Appendices:
A) Model specifications (likelihoods, priors, hyperpriors) and Gibbs updates. B) Implementation details (convergence, warmup, inference, reference alignment, label switching diagnostic). C) Additional simulation study results, including sensitivity analyses for sparse data and changing rank range. D) Additional PCWAG data application results. (AppendixA.pdf, etc.)
R package:
The R package bayesNMF is publicly available on GitHub at jennalandy/bayesNMF. Documented by README.md and vignettes.
Code:
All code needed to reproduce simulation studies and data analysis is publicly available on GitHub at jennalandy/bayesNMF_PAPER, documented by README.md.
Dataset:
The PCAWG data is available through the ICGC ARGO data portal: access instructions link. Instructions are also included in the code file 1_pcawg_data.qmd.
Appendix A Appendix A: Model Specifications
For each model, we provide a formal model definition as well as the Gibbs updates, separately for parameters and prior parameters. If applicable, MH-within-Gibbs parameter updates are also documented with targets, proposals, acceptance ratios, and MH steps. Some pieces are identical across all models, in which case the shared specification is reported in Section A.0.1 and the shared update is reported in Section A.0.2. Table A.3 summarizes the six models implemented in bayesNMF and links to the applicable subsection of this Appendix.
Throughout this appendix, let $\hat{M}$ be the reconstruction $\hat{P}\hat{A}\hat{E}$ such that $\hat{M}_{kg}=(\hat{P}\hat{E})_{kg}=\sum_{n}\hat{P}_{kn}\hat{E}_{ng}$. Similarly, let $\hat{M}_{kg/n}$ be the reconstructed value of $M_{kg}$ after excluding signature $n$: $\hat{M}_{kg/n}=\sum_{n^{\prime}\neq n}\hat{P}_{kn^{\prime}}\hat{E}_{n^{\prime}g}$.
Further, when in reference to a potential new value $P^{*}_{kn}$, let $P^{*}$ be the $P$ matrix with element $k,n$ replaced by $P^{*}_{kn}$. Similarly, when in reference to a potential new value $E^{*}_{ng}$, let $E^{*}$ be the $E$ matrix with element $n,g$ replaced by $E^{*}_{ng}$.
The value $\bar{M}$ is the mean across all count values in the data matrix $M$, representing the general scale of the data.
Likelihood Prior Sampler # Updates # Acceptance Ratios Section
Poisson TruncNormal MH-within-Gibbs $3(KN+NG)$ $KN+NG$ A.1
Poisson Exponential MH-within-Gibbs $2(KN+NG)$ $KN+NG$ A.2
Poisson Gamma Standard Gibbs $3(KN+NG)+KNG$ 0 A.3
Poisson Exponential Standard Gibbs $2(KN+NG)+KNG$ 0 A.4
Normal TruncNormal Standard Gibbs $3(KN+NG)+G$ 0 A.5
Normal Exponential Standard Gibbs $2(KN+NG)+G$ 0 A.6
Table A.3: List of models implemented in bayesNMF, reporting the number of parameter/prior parameter updates and the number of acceptance rates computed per iteration of the Gibbs sampler. Both values directly impact compute time and memory consumption. Typically $N<<K,G$. MH-within-Gibbs samplers are the new contribution of this work.
A.0.1 Definitions Shared Across Models
The prior and hyperprior structure on signature inclusion is identical across all models:
Priors: $A_{nn}\sim\text{Bernoulli}(q)$
$\text{where }q=\begin{cases}0.4/N&\text{for }R=0\\ R/N&\text{for }R=1,...,N-1\\ 1-0.4/N&\text{for }R=N\end{cases},$
Hyperpriors: $p(R=r)=1/(N+1)\text{ for }r=0,...,N.$
A.0.2 Parameter Gibbs Updates Shared Across Models
Gibbs updates for $R$ and $A$ are identical across all models, with the appropriate likelihood (Poisson or Normal) plugged in for p(M|)p(M|...):
[…]
$\text{where }q^{(r)}=\begin{cases}0.4/N&\text{if }r=0\\ r/N&\text{if }r\in\{1,...,N-1\}\\ 1-0.4/N&\text{if }r=N\end{cases}$
$A_{nn}|$... $\sim\text{Bernoulli}\left(\frac{p_{1}}{p_{1}+p_{0}}\right)$
$\text{where }p_{a}=\begin{cases}q^{a}(1-q)^{1-a}\cdot p\left(M|A_{nn}=a,...\right)^{\gamma}&\text{ in BFI}\\ q^{a}(1-q)^{1-a}\cdot\left[p\left(M|A_{nn}=a,...\right)\cdot G^{-\frac{1}{2}(K+G)N^{(a)}}\right]^{\gamma}&\text{ in SBFI}\end{cases}$
$\text{and }q=q^{(R)}$
When the latent rank is provided, the tempering parameter $\gamma$ is always 1. Otherwise, $\gamma$ starts at 0 and gradually reaches $1$ approximately 20% of the way through sampling. If $\gamma=0$, the update distribution is equivalent to the prior, and if $\gamma=1$, it is the true full conditional. No tempered samples are used for inference.
A.1 Poisson - TruncNormal + MH
A.1.1 Model Definition
Likelihood: $M_{kg}\sim\text{Poisson}\left(\sum_{n}P_{kn}A_{nn}E_{ng}\right)$
Priors: $P_{kn}\sim\text{TruncNormal}_{[0,\infty)}\left(\mu^{P}_{kn},(\sigma^{P}_{kn})^{2}\right)$
$E_{ng}\sim\text{TruncNormal}_{[0,\infty)}\left(\mu^{E}_{ng},(\sigma^{E}_{ng})^{2}\right)$
Hyperpriors: $\mu^{P}_{kn}\sim\text{Normal}\left(m^{P}_{kn}=0,(s^{2}_{kn})^{2}=\sqrt{\bar{M}/N}\right)$
$(\sigma^{P}_{kn})^{2}\sim\text{InverseGamma}\left(a^{P}_{kn}=N+1,b^{P}_{kn}=\sqrt{N}\right)$
$\mu^{E}_{ng}\sim\text{Normal}\left(m^{E}_{ng}=0,(s^{E}_{ng})^{2}=\sqrt{\bar{M}/N}\right)$
$(\sigma^{E}_{ng})^{2}\sim\text{InverseGamma}\left(a^{E}_{ng}=N+1,b^{E}_{ng}=\sqrt{N}\right)$
The prior and hyperprior structure on $A$ is defined in Section A.0.1. From here on, let $\theta$ to refer to all prior and hyperprior parameters, and assume any use of “TruncNorm” uses the bounds $0,\infty)$.
This model has $KN+NG$ parameters and $2(KN+NG)$ prior parameters to update on each iteration. MH-within-Gibbs steps are used for $P$ and $E$, meaning $KN+NG$ acceptance ratios must be computed.
A.1.2 Prior Parameter Gibbs Updates: Full Conditional Distributions
$\mu^{P}_{kn}|$... $\sim\text{Normal}\left(\mu=\frac{m^{P}_{kn}/s^{E}_{kn}+P_{kn}/(\sigma^{P}_{kn})^{2}}{1/s^{E}_{kn}+1/(\sigma^{P}_{kn})^{2}},\sigma^{2}=\frac{1}{{1/s^{E}_{kn}+1/(\sigma^{P}_{kn})^{2}}}\right)$
$(\sigma^{P}_{kn})^{2}|$... $\sim\text{InverseGamma}\left(\alpha=a^{P}_{kn}+\frac{1}{2},\beta=b^{P}_{kn}+\frac{1}{2}(P_{kn}-\mu^{P}_{kn})^{2}\right)$
$\mu^{E}_{ng}|$... $\sim\text{Normal}\left(\mu=\frac{m^{E}_{ng}/s^{E}_{ng}+E_{ng}/(\sigma^{E}_{ng})^{2}}{1/s^{E}_{ng}+1/(\sigma^{E}_{ng})^{2}},\sigma^{2}=\frac{1}{{1/s^{E}_{ng}+1/(\sigma^{E}_{ng})^{2}}}\right)$
$(\sigma^{E}_{ng})^{2}|$... $\sim\text{InverseGamma}\left(\alpha=a^{E}_{ng}+\frac{1}{2},\beta=b^{E}_{ng}+\frac{1}{2}(E_{ng}-\mu^{E}_{ng})^{2}\right)$
The update for $R$ is defined in Section A.0.2.
A.1.3 Parameter Gibbs Updates
The update for $A_{nn}$ is defined in Section A.0.2.
A.1.4 Parameter Gibbs Updates: MH-Within-Gibbs steps
Targets
$f(P^{*}_{kn}|P,E,M,\theta)$ $\propto p_{\text{TruncNorm}}\left(P^{*}_{kn}|\mu^{P}_{kn},(\sigma^{P}_{kn})^{2}\right)\prod_{g}p_{\text{Poisson}}\left(M_{kg}|\lambda=(P^{*}E)_{kg}\right)$
$f(E^{*}_{ng}|P,E,M,\theta)$ $\propto p_{\text{TruncNorm}}\left(E^{*}_{ng}|\mu^{E}_{ng},(\sigma^{E}_{ng})^{2}\right)\prod_{k}p_{\text{Poisson}}\left(M_{kg}|\lambda=(PE^{*})_{kg}\right)$
Proposals
$g(P^{*}_{kn}|P,E,M,\theta)$ $\propto p_{\text{TruncNorm}}(P^{*}_{kn}|\mu^{P}_{kn},(\sigma^{P}_{kn})^{2})\prod_{g}p_{\text{Normal}}\left(M_{kg}|\mu=(P^{*}E)_{kg},\sigma^{2}=(PE)_{kg}\right)$
$\sim\text{TruncNorm}_{[0,\infty)}\left(\mu=m^{P}_{kn}(M,P,A,E,\theta),\sigma^{2}=s^{P}_{kn}(M,P,E,\theta)\right)$
$m^{P}_{kn}(M,P,A,E,\theta)=\frac{\mu^{P}_{kn}/(\sigma^{P}_{kn})^{2}+\sum_{g=1}^{G}A_{nn}E_{ng}(M_{kg}-\hat{M}_{kg/n})/(PE)_{kg}}{1/(\sigma^{P}_{kn})^{2}+\sum_{g=1}^{G}A_{nn}E_{ng}^{2}/(PE)_{kg}}$
$s^{P}_{kn}(M,P,A,E,\theta)=\frac{1}{1/(\sigma^{P}_{kn})^{2}+\sum_{g=1}^{G}A_{nn}E_{ng}^{2}/(PE)_{kg}}$
$g(E^{*}_{ng}|P,E,M,\theta)$ $\propto p_{\text{TruncNorm}}(E^{*}_{ng}|\mu^{E}_{ng},(\sigma^{E}_{ng})^{2})\prod_{k}p_{\text{Normal}}\left(M_{kg}|\mu=(PE^{*})_{kg},\sigma^{2}=(PE)_{kg}\right)$
$\sim\text{TruncNorm}_{[0,\infty)}\left(\mu=m^{E}_{ng}(M,P,A,E,\theta),\sigma^{2}=s^{E}_{ng}(M,P,A,E,\theta)\right)$
$m^{E}_{ng}(M,P,A,E,\theta)=\frac{\mu^{E}_{ng}/(\sigma^{E}_{ng})^{2}+\sum_{k=1}^{K}P_{kn}A_{nn}(M_{kg}-\hat{M}_{kg/n})/(PE)_{kg}}{1/(\sigma^{E}_{ng})^{2}+\sum_{k=1}^{K}P_{kn}^{2}A_{nn}/(PE)_{kg}}$
$s^{E}_{ng}(M,P,A,E,\theta)=\frac{1}{1/(\sigma^{E}_{ng})^{2}+\sum_{k=1}^{K}P_{kn}^{2}A_{nn}/(PE)_{kg}}$
Acceptance Ratios
$a^{P}_{kn}$ $=\frac{f(P^{*}_{kn}|P_{-kn},E,M,\theta)g(P_{kn}|P^{*},E,M,\theta)}{f(P_{kn}|P_{-kn},E,M,\theta)g(P^{*}_{kn}|P,E,M,\theta)}$
$=\prod_{g}\frac{p_{\text{Poisson}}\left(M_{kg}|\lambda=(P^{*}E)_{kg}\right)\cdot p_{\text{Normal}}\left(M_{kg}|\mu=(PE)_{kg},\sigma^{2}=(P^{*}E)_{kg}\right)}{p_{\text{Poisson}}\left(M_{kg}|\lambda=(PE)_{kg}\right)\cdot p_{\text{Normal}}\left(M_{kg}|\mu=(P^{*}E)_{kg},\sigma^{2}=(PE)_{kg}\right)}$
$a^{E}_{ng}$ $=\frac{f(E^{*}_{ng}|P,E_{-ng},M,\theta)g(E_{ng}|P,E^{*},M,\theta)}{f(E_{ng}|P,E_{-ng},M,\theta)g(E^{*}_{ng}|P,E,M,\theta)}$
$=\prod_{g}\frac{p_{\text{Poisson}}\left(M_{kg}|\lambda=(PE^{*})_{kg}\right)\cdot p_{\text{Normal}}\left(M_{kg}|\mu=(PE)_{kg},\sigma^{2}=(PE^{*})_{kg}\right)}{p_{\text{Poisson}}\left(M_{kg}|\lambda=(PE)_{kg}\right)\cdot p_{\text{Normal}}\left(M_{kg}|\mu=(PE^{*})_{kg},\sigma^{2}=(PE)_{kg}\right)}$
Updates / MH steps
[…]
A.2 Poisson - Exponential + MH
A.2.1 Model Definition
Likelihood: $M_{kg}\sim\text{Poisson}\left(\sum_{n}P_{kn}A_{nn}E_{ng}\right)$
Priors: $P_{kn}\sim\text{Exponential}\left(\lambda^{P}_{kn}\right)$
$E_{ng}\sim\text{Exponential}\left(\lambda^{E}_{ng}\right)$
Hyperpriors: $\lambda^{P}_{kn}\sim\text{Gamma}\left(a^{P}_{kn}=10\sqrt{N},b^{P}_{kn}=10\sqrt{\bar{M}}\right)$
$\lambda^{E}_{ng}\sim\text{Gamma}\left(a^{E}_{ng}=10\sqrt{N},b^{E}_{ng}=10\sqrt{\bar{M}}\right)$
The prior and hyperprior structure on $A$ is defined in Section A.0.1. From here on, let $\theta$ to refer to all prior and hyperprior parameters.
This model has $KN+NG$ parameters and $KN+NG$ prior parameters to update on each iteration. MH-within-Gibbs steps are used for $P$ and $E$, meaning $KN+NG$ acceptance ratios must be computed.
A.2.2 Prior Parameter Full Conditional Distributions
$\lambda^{P}_{kn}|$... $\sim\text{Gamma}\left(a^{P}_{kn}+1,b^{P}_{kn}+P_{kn}\right)$
$\lambda^{E}_{ng}|$... $\sim\text{Gamma}\left(a^{E}_{ng}+1,b^{E}_{ng}+E_{ng}\right)$
The update for $R$ is defined in Section A.0.2.
A.2.3 Parameter Gibbs Updates
The update for $A_{nn}$ is defined in Section A.0.2.
A.2.4 Parameter Gibbs Updates: MH-Within-Gibbs steps
Targets
$f(P^{*}_{kn}|P,E,M,\theta)$ $\propto p_{\text{Exp}}\left(P^{*}_{kn}|\lambda^{P}_{kn}\right)\prod_{g}p_{\text{Poisson}}\left(M_{kg}|\lambda=(P^{*}E)_{kg}\right)$
$f(E^{*}_{ng}|P,E,M,\theta)$ $\propto p_{\text{Exp}}\left(E^{*}_{ng}|\lambda^{E}_{ng}\right)\prod_{k}p_{\text{Poisson}}\left(M_{kg}|\lambda=(PE^{*})_{kg}\right)$
Proposals
$g(P^{*}_{kn}|P,E,M,\theta)$ $\propto p_{\text{Exp}}(P^{*}_{kn}|\lambda^{P}_{kn})\prod_{g}p_{\text{Normal}}\left(M_{kg}|\mu=(P^{*}E)_{kg},\sigma^{2}=(PE)_{kg}\right)$
$\sim\text{TruncNorm}_{[0,\infty)}\left(\mu=m^{P}_{kn}(M,P,A,E,\theta),\sigma^{2}=s^{P}_{kn}(M,P,E,\theta)\right)$
$m^{P}_{kn}(M,P,A,E,\theta)=\frac{\sum_{g=1}^{G}A_{nn}E_{ng}(M_{kg}-\hat{M}_{kg/n})/(PE)_{kg}-\lambda^{P}_{kn}}{\sum_{g=1}^{G}A_{nn}E_{ng}^{2}/(PE)_{kg}}$
$s^{P}_{kn}(M,P,A,E,\theta)=\frac{1}{\sum_{g=1}^{G}A_{nn}E_{ng}^{2}/(PE)_{kg}}$
$g(E^{*}_{ng}|P,E,M,\theta)$ $\propto p_{\text{Exp}}(E^{*}_{ng}|\lambda^{E}_{ng})\prod_{k}p_{\text{Normal}}\left(M_{kg}|\mu=(PE^{*})_{kg},\sigma^{2}=(PE)_{kg}\right)$
$\sim\text{TruncNorm}_{[0,\infty)}\left(\mu=m^{E}_{ng}(M,P,A,E,\theta),\sigma^{2}=s^{E}_{ng}(M,P,A,E,\theta)\right)$
$m^{E}_{ng}(M,P,A,E,\theta)=\frac{\sum_{k=1}^{K}P_{kn}A_{nn}(M_{kg}-\hat{M}_{kg/n})/(PE)_{kg}-\lambda^{E}_{ng}}{\sum_{k=1}^{K}P_{kn}^{2}A_{nn}/(PE)_{kg}}$
$s^{E}_{ng}(M,P,A,E,\theta)=\frac{1}{\sum_{k=1}^{K}P_{kn}^{2}A_{nn}/(PE)_{kg}}$
Acceptance Ratios
$a^{P}_{kn}$ $=\frac{f(P^{*}_{kn}|P_{-kn},E,M,\theta)g(P_{kn}|P^{*},E,M,\theta)}{f(P_{kn}|P_{-kn},E,M,\theta)g(P^{*}_{kn}|P,E,M,\theta)}$
$=\prod_{g}\frac{p_{\text{Poisson}}\left(M_{kg}|\lambda=(P^{*}E)_{kg}\right)\cdot p_{\text{Normal}}\left(M_{kg}|\mu=(PE)_{kg},\sigma^{2}=(P^{*}E)_{kg}\right)}{p_{\text{Poisson}}\left(M_{kg}|\lambda=(PE)_{kg}\right)\cdot p_{\text{Normal}}\left(M_{kg}|\mu=(P^{*}E)_{kg},\sigma^{2}=(PE)_{kg}\right)}$
$a^{E}_{ng}$ $=\frac{f(E^{*}_{ng}|P,E_{-ng},M,\theta)g(E_{ng}|P,E^{*},M,\theta)}{f(E_{ng}|P,E_{-ng},M,\theta)g(E^{*}_{ng}|P,E,M,\theta)}$
$=\prod_{g}\frac{p_{\text{Poisson}}\left(M_{kg}|\lambda=(PE^{*})_{kg}\right)\cdot p_{\text{Normal}}\left(M_{kg}|\mu=(PE)_{kg},\sigma^{2}=(PE^{*})_{kg}\right)}{p_{\text{Poisson}}\left(M_{kg}|\lambda=(PE)_{kg}\right)\cdot p_{\text{Normal}}\left(M_{kg}|\mu=(PE^{*})_{kg},\sigma^{2}=(PE)_{kg}\right)}$
Updates / MH steps
[…]
A.3 Poisson - Gamma
A.3.1 Model Definition
Likelihood: $M_{kg}=\sum_{n}Z_{kng}$
$Z_{kng}\sim\text{Poisson}\left(P_{kn}A_{nn}E_{ng}\right)$
Priors: $P_{kn}\sim\text{Gamma}\left(\alpha^{P}_{kn},\beta^{P}_{kn}\right)$
$E_{ng}\sim\text{Gamma}\left(\alpha^{E}_{ng},\beta^{E}_{ng}\right)$
Hyperpriors: $\beta^{P}_{kn}\sim\text{Gamma}(a^{P}_{kn}=10\sqrt{N},b^{P}_{kn}=10)$
$\alpha^{P}_{kn}\sim\text{Gamma}(c^{P}_{kn}=10\sqrt{\bar{M}},d^{P}_{kn}=10)$
$\beta^{E}_{ng}\sim\text{Gamma}(a^{E}_{ng}=10\sqrt{N},b^{E}_{ng}=10)$
$\alpha^{E}_{ng}\sim\text{Gamma}(c^{E}_{ng}=10\sqrt{\bar{M}},d^{E}_{ng}=10)$
The prior and hyperprior structure on $A$ is defined in Section A.0.1. From here on, let $\theta$ to refer to all prior and hyperprior parameters.
This model has $KN+NG+KNG$ parameters and $2(KN+NG)$ prior parameters to update on each iteration.
A.3.2 Prior Parameter Gibbs Updates: Full Conditional Distributions
$\beta^{P}_{kn}$ $\sim\text{Gamma}(a^{P}_{kn}+\alpha^{P}_{kn},b^{P}_{kn}+P_{kn})$
[…]
$\beta^{E}_{ng}$ $\sim\text{Gamma}(a^{E}_{ng}+\alpha^{E}_{ng},b^{E}_{ng}+E_{ng})$
[…]
$\alpha^{P}_{kn}$ and $\alpha^{E}_{ng}$ are sampled using Adaptive Rejection Metropolis Sampling (ARMS) as implemented by the armspp R software package [Bertolacci, 2019]. The update for $R$ is defined in Section A.0.2.
A.3.3 Parameter Gibbs Updates
$P_{kn}|\dots$ $\sim\text{Gamma}\left(\alpha^{P}_{kn}+\sum_{g}Z_{kng},\beta^{P}_{kn}+A_{nn}\sum_{g}E_{ng}\right)$
$E_{ng}|\dots$ $\sim\text{Gamma}\left(\alpha^{E}_{ng}+\sum_{k}Z_{kng},\beta^{E}_{ng}+A_{nn}\sum_{k}P_{kn}\right)$
$\vec{Z}_{k\cdot g}|\dots$ $\sim\text{Multinomial}\left(P_{k1}A_{11}E_{1g},\dots,P_{kN}A_{NN}E_{Ng}\right)$
The update for $A_{nn}$ is defined in Section A.0.2.
A.4 Poisson - Exponential
A.4.1 Model Definition
Likelihood: $M_{kg}=\sum_{n}Z_{kng}$
$Z_{kng}\sim\text{Poisson}\left(P_{kn}A_{nn}E_{ng}\right)$
Priors: $P_{kn}\sim\text{Exponential}\left(\lambda^{P}_{kn}\right)$
$E_{ng}\sim\text{Exponential}\left(\lambda^{E}_{ng}\right)$
Hyperpriors: $\lambda^{P}_{kn}\sim\text{Gamma}\left(a^{P}_{kn}=10\sqrt{N},b^{P}_{kn}=10\sqrt{\bar{M}}\right)$
$\lambda^{E}_{ng}\sim\text{Gamma}\left(a^{E}_{ng}=10\sqrt{N},b^{E}_{ng}=10\sqrt{\bar{M}}\right)$
The prior and hyperprior structure on $A$ is defined in Section A.0.1. From here on, let $\theta$ to refer to all prior and hyperprior parameters.
This model has $KN+NG+KNG$ parameters and $KN+NG$ prior parameters to update on each iteration.
A.4.2 Prior Parameter Gibbs Updates: Full Conditional Distributions
$\lambda^{P}_{kn}|$... $\sim\text{Gamma}\left(a^{P}_{kn}+1,b^{P}_{kn}+P_{kn}\right)$
$\lambda^{E}_{ng}|$... $\sim\text{Gamma}\left(a^{E}_{ng}+1,b^{E}_{ng}+E_{ng}\right)$
The update for $R$ is defined in Section A.0.2.
A.4.3 Parameter Gibbs Updates
$P_{kn}|\dots$ $\sim\text{Gamma}\left(1+\sum_{g}Z_{kng},\lambda^{P}_{kn}+A_{nn}\sum_{g}E_{ng}\right)$
$E_{ng}|\dots$ $\sim\text{Gamma}\left(1+\sum_{k}Z_{kng},\lambda^{E}_{ng}+A_{nn}\sum_{k}P_{kn}\right)$
$\vec{Z}_{k\cdot g}|\dots$ $\sim\text{Multinomial}\left(P_{k1}A_{11}E_{1g},\dots,P_{kN}A_{NN}E_{Ng}\right)$
The update for $A_{nn}$ is defined in Section A.0.2.
A.5 Normal - TruncNormal
A.5.1 Model Definition
Likelihood: $M_{kg}\sim\text{Normal}\left(\sum_{n}P_{kn}A_{nn}E_{ng},\sigma^{2}_{g}\right)$
Priors: $P_{kn}\sim\text{TruncNormal}_{[0,\infty)}\left(\mu^{P}_{kn},(\sigma^{P}_{kn})^{2}\right)$
$E_{ng}\sim\text{TruncNormal}_{[0,\infty)}\left(\mu^{E}_{ng},(\sigma^{E}_{ng})^{2}\right)$
$\sigma^{2}_{g}\sim\text{InverseGamma}(\alpha_{g}=3,\beta_{g}=3)$
Hyperpriors: $\mu^{P}_{kn}\sim\text{Normal}\left(m^{P}_{kn}=0,(s^{2}_{kn})^{2}=\sqrt{\bar{M}/N}\right)$
$(\sigma^{P}_{kn})^{2}\sim\text{InverseGamma}\left(a^{P}_{kn}=N+1,b^{P}_{kn}=\sqrt{N}\right)$
$\mu^{E}_{ng}\sim\text{Normal}\left(m^{E}_{ng}=0,(s^{E}_{ng})^{2}=\sqrt{\bar{M}/N}\right)$
$(\sigma^{E}_{ng})^{2}\sim\text{InverseGamma}\left(a^{E}_{ng}=N+1,b^{E}_{ng}=\sqrt{N}\right)$
The prior and hyperprior structure on $A$ is defined in Section A.0.1. From here on, let $\theta$ to refer to all prior and hyperprior parameters, and assume any use of “TruncNorm” uses the bounds $0,\infty)$.
This model has $KN+NG+G$ parameters and $2(KN+NG)$ prior parameters to update on each iteration.
A.5.2 Prior Parameter Gibbs Updates: Full Conditional Distributions
$\mu^{P}_{kn}|$... $\sim\text{Normal}\left(\mu=\frac{m^{P}_{kn}/s^{E}_{kn}+P_{kn}/(\sigma^{P}_{kn})^{2}}{1/s^{E}_{kn}+1/(\sigma^{P}_{kn})^{2}},\sigma^{2}=\frac{1}{{1/s^{E}_{kn}+1/(\sigma^{P}_{kn})^{2}}}\right)$
$(\sigma^{P}_{kn})^{2}|$... $\sim\text{InverseGamma}\left(\alpha=a^{P}_{kn}+\frac{1}{2},\beta=b^{P}_{kn}+\frac{1}{2}(P_{kn}-\mu^{P}_{kn})^{2}\right)$
$\mu^{E}_{ng}|$... $\sim\text{Normal}\left(\mu=\frac{m^{E}_{ng}/s^{E}_{ng}+E_{ng}/(\sigma^{E}_{ng})^{2}}{1/s^{E}_{ng}+1/(\sigma^{E}_{ng})^{2}},\sigma^{2}=\frac{1}{{1/s^{E}_{ng}+1/(\sigma^{E}_{ng})^{2}}}\right)$
$(\sigma^{E}_{ng})^{2}|$... $\sim\text{InverseGamma}\left(\alpha=a^{E}_{ng}+\frac{1}{2},\beta=b^{E}_{ng}+\frac{1}{2}(E_{ng}-\mu^{E}_{ng})^{2}\right)$
The update for $R$ is defined in Section A.0.2.
A.5.3 Parameter Gibbs Updates
$P_{kn}|$... $\sim\text{TruncNorm}_{[0,\infty)}\left(\mu=m^{P}_{kn}(M,A,E,\theta),\sigma^{2}=s^{P}_{kn}(M,A,E,\theta)\right)$
$m^{P}_{kn}(M,A,E,\theta)=\frac{\mu^{P}_{kn}/(\sigma^{P}_{kn})^{2}+\sum_{g=1}^{G}A_{nn}E_{ng}(M_{kg}-\hat{M}_{kg/n})/\sigma^{2}_{g}}{1/(\sigma^{P}_{kn})^{2}+\sum_{g=1}^{G}A_{nn}E_{ng}^{2}/\sigma^{2}_{g}}$
$s^{P}_{kn}(M,A,E,\theta)=\frac{1}{1/(\sigma^{P}_{kn})^{2}+\sum_{g=1}^{G}A_{nn}E_{ng}^{2}/\sigma^{2}_{g}}$
$E_{ng}|$... $\sim\text{TruncNorm}_{[0,\infty)}\left(\mu=m^{E}_{ng}(M,P,A,\theta),\sigma^{2}=s^{E}_{ng}(M,P,A,\theta)\right)$
$m^{E}_{ng}(M,P,A,\theta)=\frac{\mu^{E}_{ng}/(\sigma^{E}_{ng})^{2}+\sum_{k=1}^{K}P_{kn}A_{nn}(M_{kg}-\hat{M}_{kg/n})/\sigma^{2}_{g}}{1/(\sigma^{E}_{ng})^{2}+\sum_{k=1}^{K}P_{kn}^{2}A_{nn}/\sigma^{2}_{g}}$
$s^{E}_{ng}(M,P,A,\theta)=\frac{1}{1/(\sigma^{E}_{ng})^{2}+\sum_{k=1}^{K}P_{kn}^{2}A_{nn}/\sigma^{2}_{g}}$
$\sigma^{2}_{g}|$... $\sim\text{InverseGamma}\left(\alpha_{g}+K/2,\beta_{g}+\sum_{k}(M_{kg}-\hat{M}_{kg})^{2}/2\right)$
The update for $A_{nn}$ is defined in Section A.0.2.
A.6 Normal - Exponential
A.6.1 Model Definition
Likelihood: $M_{kg}\sim\text{Normal}\left(\sum_{n}P_{kn}A_{nn}E_{ng},\sigma^{2}_{g}\right)$
Priors: $P_{kn}\sim\text{Exponential}\left(\lambda^{P}_{kn}\right)$
$E_{ng}\sim\text{Exponential}\left(\lambda^{E}_{ng}\right)$
$\sigma^{2}_{g}\sim\text{InverseGamma}\left(\alpha_{g}=3,\beta_{g}=3\right)$
Hyperpriors: $\lambda^{P}_{kn}\sim\text{Gamma}\left(a^{P}_{kn}=10\sqrt{N},b^{P}_{kn}=10\sqrt{\bar{M}}\right)$
$\lambda^{E}_{ng}\sim\text{Gamma}\left(a^{E}_{ng}=10\sqrt{N},b^{E}_{ng}=10\sqrt{\bar{M}}\right)$
The prior and hyperprior structure on $A$ is defined in Section A.0.1. From here on, let $\theta$ to refer to all prior and hyperprior parameters.
This model has $KN+NG+G$ parameters and $KN+NG$ prior parameters to update on each iteration.
A.6.2 Prior Parameter Full Conditional Distributions
$\lambda^{P}_{kn}|$... $\sim\text{Gamma}\left(a^{P}_{kn}+1,b^{P}_{kn}+P_{kn}\right)$
$\lambda^{E}_{ng}|$... $\sim\text{Gamma}\left(a^{E}_{ng}+1,b^{E}_{ng}+E_{ng}\right)$
The update for $R$ is defined in Section A.0.2.
A.6.3 Parameter Gibbs Updates
$P_{kn}$... $\sim\text{TruncNorm}_{[0,\infty)}\left(\mu=m^{P}_{kn}(M,A,E,\theta),\sigma^{2}=s^{P}_{kn}(M,A,E,\theta)\right)$
$m^{P}_{kn}(M,A,E,\theta)=\frac{\sum_{g=1}^{G}A_{nn}E_{ng}(M_{kg}-\hat{M}_{kg/n})/\sigma^{2}_{g}-\lambda^{P}_{kn}}{\sum_{g=1}^{G}A_{nn}E_{ng}^{2}/\sigma^{2}_{g}}$
$s^{P}_{kn}(M,A,E,\theta)=\frac{1}{\sum_{g=1}^{G}A_{nn}E_{ng}^{2}/\sigma^{2}_{g}}$
$E_{ng}|$... $\sim\text{TruncNorm}_{[0,\infty)}\left(\mu=m^{E}_{ng}(M,P,A,\theta),\sigma^{2}=s^{E}_{ng}(M,P,A,\theta)\right)$
$m^{E}_{ng}(M,P,A,\theta)=\frac{\sum_{k=1}^{K}P_{kn}A_{nn}(M_{kg}-\hat{M}_{kg/n})/\sigma^{2}_{g}-\lambda^{E}_{ng}}{\sum_{k=1}^{K}P_{kn}^{2}A_{nn}/\sigma^{2}_{g}}$
$s^{E}_{ng}(M,P,A,\theta)=\frac{1}{\sum_{k=1}^{K}P_{kn}^{2}A_{nn}/\sigma^{2}_{g}}$
$\sigma^{2}_{g}|$... $\sim\text{InverseGamma}\left(\alpha_{g}+K/2,\beta_{g}+\sum_{k}(M_{kg}-\hat{M}_{kg})^{2}/2\right)$
The update for $A_{nn}$ is defined in Section A.0.2.
Appendix B Appendix B: Implementation Details
B.1 Inference
Inference is performed on the last 1000 posterior samples. No tempered or non-MH samples are used for final posterior inference. This same strategy is on periodic windows of samples to check for convergence.
When learning rank, we first compute $\hat{A}$, the maximum a-posteriori (MAP) estimate of the signature inclusion matrix, at its mode. We then subset inference samples to those that match this MAP.
To summarize the inference samples, we next remove the scale indeterminacy of NMF by normalizing each iteration’s $P^{(i)}$ and $E^{(i)}$ matrices for columns of $P^{(i)}$ to sum to one:
$E^{\prime}_{ng}=E_{ng}*\sum_{k}P_{kn}$
$P^{\prime}_{kn}=P_{kn}/\sum_{k^{\prime\prime}}P_{k^{\prime\prime}n}.$
This rescaling ensures the reconstruction has not changed, $PE=P^{\prime}E^{\prime}$.
Then, final estimates $\hat{P}$ and $\hat{E}$ and 95% credible intervals are element-wise means and 2.5th and 97.5th percentiles, respectively. The estimated reconstruction $\hat{M}$ is the product $\hat{P}\hat{A}\hat{E}$.
B.2 Convergence
[…]
We instead define a programmatic convergence criterion that is appropriate for any scale of data and ensures that the various models reach the same level of convergence. This convergence criterion is defined on a running-metrics plot, built off of the concept of a running means plot. We assume that if posterior samples have converged, then the average over windows of fixed lengths should not change as the window shifts, and thus a scalar metric of this average should not change either.
Each time convergence is checked, inference over the window follows Section B.1. Metrics RMSE and KL-Divergence are computed on the reconstruction $\hat{M}=\hat{P}\hat{A}\hat{E}$. However, scale-dependent metrics of log-likelihood, BIC, and log-posterior don’t make sense to compute on the rescaled $\hat{P}$ and $\hat{E}$. Instead, these window metrics are the average over per-sample metrics within that window.
All parameters can be adjusted as input to the new_convergence_controls function. Default settings used for all results in our paper consider the convergence of log posterior (metric) with a window size of $1000$ samples (MAP_over) and window shifts of $100$ samples (MAP_every). The package also allows for convergence of BIC, log likelihood, RMSE, or KL-divergence. A change of less than 0.10.1% in log posterior is considered “no change” (tol). We only check for convergence once the sampler has run for at least a minimum number of iterations (miniters) and none of the samples in the window were tempered. We say a sampler has converged if the log posterior has not changed in the past 5 status checks (Ninarow_nochange), has not improved in the past 10 status checks (Ninarow_nobest), or has reached the maximum number of iterations specified (maxiters).
Figure B.1 shows a running metrics plot using the default convergence control parameters on an example simulated dataset with fixed rank. Note the extremely tight y-axis for log-posterior distribution (was NA at iteration 1000).
Figure B.1: Running metrics plot for an example simulated dataset with fixed rank. Vertical blue line indicates convergence (see Section B.2) and thus the switch from “accept all” to true MH samples (see Section B.3). The highlighted light blue rectangle indicates samples used for final inference.
B.3 Warmup and MH Steps
We find in practice that the MH-within-Gibbs steps are slow to explore when starting from prior samples. Instead, we accept all proposals in a “warm-up” phase to find the high-posterior space shared by the Normal and Poisson Bayesian NMFs. Once the sampler has converged, it switches to true MH steps for the last 2000 iterations, using only the last 1000 for inference.
We see in practice that after switching to true MH steps, the RMSE increases slightly and KL-Divergence decreases slightly (Figure B.1). This makes sense as we move away from samples closer to a Normal Bayesian NMF posterior space (whose likelihood optimizes RMSE) to a Poisson Bayesian NMF posterior space (whose likelihood optimizes KL-Divergence).
B.4 Alignment to Reference Signatures
In order to interpret and evaluate results, we frequently wish to compare estimated factors (columns of $\hat{P}$) to reference factors (columns of $P$) from previous analysis or domain knowledge, commonly from the COSMIC database [Consortium, 2020] in the case of mutational signatures. We use the following approach to align estimated signatures to ground truth in our simulation studies as well.
We follow existing mutational signatures works [Rosales et al., 2016, Gori and Baez-Ortega, 2018, Alexandrov et al., 2020, Grabski et al., 2023b] and focus on the scale-independent measure of cosine similarities between signatures. To align estimated factors to the reference factors, we first create a similarity matrix $S$, where each element is the cosine similarity between a column of $P$ and a column of $\hat{P}$, $S_{ij}=cos(P_{i},\hat{P}_{j})$. The Hungarian algorithm [Kuhn, 1955] is used on $1\cdot S$, which solves the combinatorial optimization problem to reorder rows and columns to minimize the sum of values on the diagonal. The resulting aligned similarity matrix, $S^{*}$, has an optimal alignment in terms of total aligned cosine similarity. This approach only allows a single estimated signature to be aligned to each reference signature, and vice versa.
We incorporate posterior uncertainty into this process by performing factor alignment for each of the posterior samples used to compute $\hat{P}$. We determine final alignments through majority voting with cosine similarity as voting weights.
B.5 Label Switching Diagnostic
Label switching is always a concern in Bayesian mixture models. This is partially mitigated by our use of hyperpriors, which avoids factors fully resetting to the prior when they are excluded. However, even with hyperpriors, label switching is still possible. To account for this, we have added additional visualization capabilities to our R software package to help diagnose label switching.
Given a set of reference factors (e.g., COSMIC mutational signatures), we visualize the closest match of each signature across iterations. If label switching occurs, the color of a latent factor will switch part way through the chain.
In our simulation studies, we see evidence of label switching only very early on in the warm-up phase (first 500 iterations or less), but labels stabilize far before samples are used for inference. For example, in Figure B.2B, There is some label switching with SBS40, which is initially aligned to estimated signature 14, then estimated signature 10, and eventually estimated signature 6 within the first few hundred iterations. SBS40 is then consistently assigned to estimated signature 6 for the remainder of the sampler’s run.
Although these individual observations do not constitute a full benchmarking, this functionality provides a diagnostic for individual use cases.
Figure B.2: Label switching diagnostic plot on simulated data with A. N = 5, B. N = 10, and C. N = 15 signatures.
Appendix C Appendix C: Additional Simulation Results
C.1 Fixed Rank Simulation Additional Results
Figure C.1 is an extended version of main text Figure 2, reporting metrics for all six models. Warm colors are Poisson-Exponential models with and without MH steps. We expect these to be near identical in minimum cosine similarity, RMSE, and KL-Divergence. Cool colors are Poisson-Gamma (with augmentation) and Poisson-Truncated Normal (with MH steps). We expect these to be close, but not identical because of the different priors used. In black and grey are Normal-likelihood models.
In addition to metrics discussed in the main text, we report mean acceptance rates and effective sample sizes. Mean acceptance rates are only reported for Poisson+MH models, and report the average acceptance ratio over all elements of the P or E matrix, averaged again over all 1000 posterior samples. The results show that acceptance rates are very high ($\geq 0.9$, on average), consistent with expectations for high-overlap, geometry-informed proposals.
Effective sample size (ESS), or the approximate number of independent samples that would correspond to a dependent chain of values $\mathbf{x}$, is estimated as follows:
$\widehat{ESS}(\mathbf{x})$ $=\frac{n}{1+2\sum_{k=1}^{K}\rho(k)}$
$\rho(k)$ $=\text{cor}\left(\mathbf{x}_{1:(n-k)},\mathbf{x}_{k:n}\right)$
$K$ $=\arg\min_{k}\{k:\rho(k)<0\}$
where $n$ is the length of vector $\mathbf{x}$, $\rho(k)$ is the correlation between $\mathbf{x}$ and itself at a lag of $k$, and $K$ is the smallest value of $k\in 1,...,n$ that results in a negative lag correlation $\rho(k)$.
Figure C.1: Metrics for all models on simulated data with fixed rank, including average acceptance rate across all elements of the P and E matrices for Poisson+MH models, and average effective sample size (ESS) across all elements of the P and E matrics (out of 1000 posterior samples).
Effective sample sizes tends to decrease with dimensionality. This is expected, as conditional dependencies tighten with more parameters (see Betancourt [2017] for an intuitive explanation). We also see a pattern of higher ESS for our Poisson+MH models than their corresponding standard Poisson samplers, indicating a more efficient exploration of the posterior space.
C.2 Comparing Point Estimates: Additional Details
Comparison Prior # datasets with # datasets with
P min cosine sim <<0.8 E min cosine sim <<0.8
Normal vs Poisson Exponential 5 2
Normal vs Poisson Gamma / TN 5 1
Poisson vs Poisson+MH Exponential 6 2
Poisson vs Poisson+MH Gamma / TN 3 0
Poisson vs Poisson Exponential 7 3
Poisson vs Poisson Gamma / TN 1 1
Table C.4: Number of points excluded from Figure 1A by comparison and prior, out of 100 datasets.
Table C.4 shows the number of points excluded from main text Figure 1A. Importantly, the number is comparable across all comparisons, even the null case comparing two chains of the standard Poisson sampler. This indicates that such deviation is reasonably expected due to the randomness of Gibbs samplers, and does not change the results reported in the main text.
C.3 Comparing Posterior Uncertainty
Figure C.2: Posterior uncertainty agreement between Poisson Bayesian NMF models in exposures matrix (left) and signatures matrix (right). Plotted are 10th (negative) and 90th (positive) percentiles of the difference in credible interval widths with standard Poisson as reference. Transparent boxes represent the null case comparing two chains of a standard Poisson sampler. Colored boxes compare Poisson + MH to standard Poisson.
C.3.1 Methods
Each sampler S returns point estimates as well as 95% credible interval bounds for elements of P, $l^{S}(P_{kn}),u^{S}(P_{kn})]$, and E, $l^{S}(E_{ng}),u^{S}(E_{ng})]$.
To compare two samplers S and Q on a given simulated dataset, we first compute credible interval widths $w$ for each element of $P$ and $E$ from each sampler. For example, for element $P_{kn}$ ,:
$w^{S}(P_{kn})$ $=u^{S}(P_{kn})-l^{S}(P_{kn}),\quad w^{Q}(P_{kn})=u^{Q}(P_{kn})-l^{Q}(P_{kn}).$ (9)
Next, we compute element-wise differences in widths between samplers, $d$:
$d^{SQ}(P_{kn})$ $=w^{S}(P_{kn})-w^{Q}(P_{kn}).$ (10)
Letting $d^{SQ}(P)$ be the flattened vector of $d^{SQ}(P_{kn})$ for all $k,n$, and similarly for $d^{SQ}(E)$, we compute the 10th and 90th percentiles of these differences, separately for $P$ and $E$. Then, comparing samplers S and Q on the given dataset can be summarized by two ranges:
$[\text{perc}_{10}(d^{SQ}(P)),\text{perc}_{90}(d^{SQ}(P))]$ (11)
$[\text{perc}_{10}(d^{SQ}(E)),\text{perc}_{90}(d^{SQ}(E))]$ (12)
Treating sampler S as a reference, these can be interpreted as reasonable ranges of how conservative (wider, lower limits) or anticonservative (narrower, upper limits) sampler Q is expected to be compared to sampler S. Narrow ranges around zero indicate a similar degree of posterior uncertainty between samplers—this, combined with near-identical point point estimates, indicates the two samplers explore the same posterior space.
C.3.2 Simulation Studies
In our simulation studies with known rank (see main body Section 5.3), we compare posterior uncertainties as described above between the Poisson + MH sampler and the standard Poisson sampler. We also compare two chains of the standard Poisson sampler to describe a null case, or the range of differences expected due to the randomness of samplers.
For each comparison $S,Q$, Figure C.2 plots the ranges $\text{perc}_{10}(d^{SQ}(P)),\text{perc}_{90}(d^{SQ}(P))]$ (left) and $\text{perc}_{10}(d^{SQ}(E)),\text{perc}_{90}(d^{SQ}(E))]$ (right) across the 100 datasets. Transparent boxes represent the null case. Instances where the colored boxes fall below the transparent boxes indicate that our Poisson + MH model is more conservative than expected.
We see that Poisson + MH samplers are generally a bit more conservative than standard Poisson samplers, and are only slightly anticonservative in the simplest cases when latent rank is 2. We also see that when the prior is an exact match (Exponential, orange), the widths are much closer to the expected variation. This is reasonable because the posterior space should change when the prior changes from a Gamma to a Truncated Normal (blue).
Overall, the scale of the y-axes indicate differences are minute, with all signature widths within 0.004 (on the probability scale) and all exposures within 2% of total mutations per sample. The difference in mutations ranges from 20 mutations per sample (seen for individuals with 2 signatures and thus an expected 2,000 mutations) to 200 mutations per sample (seen for individuals with 16 signatures and thus an expected 16,000 mutations).
The slightly wider credible intervals may in fact be a result of more efficient posterior exploration as seen with higher effective sample sizes for Poisson+MH in Section C.1. This suggests these wider credible intervals may be more accurate, while standard Poisson samplers may be too narrow.
C.4 Learning Rank Simulation Additional Results
Figure C.3: Rank bias, precision, sensitivity, and time comparison of rank learning approaches: bayesNMF with SBFI (upright triangle), and Heuristic (square), as well as SignatureAnalyzer+ARD (inverted triangle). Plotted are large dark points for medians overlapping lighter boxplots and jittered outliers. Split into N = 1-10 (top) and 11-20 (bottom).
Figure C.3 is an extended version of main text Figure 3, reporting metrics for all ranks between 1 and 20. Results match what is reported in main text 5.3.
C.5 Sensitivity Analysis: Sparse Data
Because the Normal approximation to Poisson breaks down with sparse or low counts, we investigate how bayesNMF performs on this type of data.
C.5.1 Fixed Rank
Data are simulated as in main text Section 5.2 for N = 4 signatures only and a reduced expected number of mutations per sample per signature of 100. This implies 400 mutations per sample spread across 96 mutation types, resulting in 20-40% sparsity (Figure C.4A) and low median mutational counts below 3 (Figure C.4B). This analysis only compares Poisson-Exponential and Poisson-Exponential + MH (focusing on the case of exact model alignment).
The average acceptance rate has decreased from around 0.9-0.95 in the primary simulation study to 0.85-0.9 in this sensitivity analysis (Figure C.4C). This is expected because the geometry of the Normal NMF is not as close to the Poisson NMF when counts are low. However, these acceptance rates are still relatively high, indicating our proposals are still high-overlap and geometry-informed.
With these results, we recreated main text Figure 2 comparing point estimates, which shows that even with sparse counts, Poisson+MH reaches the same point estimate as a standard Poisson sampler (Figure C.4D). We’ve also recreated Appendix Figure C.2, which shows that the Poisson+MH reaches the same posterior uncertainty as a standard Poisson sampler (Figure C.4E).
Together, these results show that even with sparse data, the Poisson-Exponential+MH model is still able to match the standard Poisson-Exponential model, both in terms of point estimates and posterior uncertainty.
Figure C.4: Data composition and results of bayesNMF with Exponential priors using fixed rank. A. Sparsity (proportion 0s) of simulated data. B. Median of mutational counts matrix of simulated datasets. C. Mean acceptance rates within each matrix across posterior samples used for inference. D. Agreement of point estimates between Poisson and Poisson + MH and the null case between repeated Poisson chains. E. Agreement of 95% credible intervals in terms of 10 and 90th percentile of width difference (see Section C.2).
C.5.2 Learning Rank
This sensitivity analysis compares Poisson+MH SBFI to SignatureAnalyzer with Truncated Normal/L2 priors.
Figure C.5: Data composition and results for learning rank with sparse data. A. Sparsity (proportion 0s) of simulated data. B. Median of mutational counts matrix of simulated datasets. C. Rank bias, precision, sensitivity, and time requirement (log scale) bayesNMF with Truncated Normal priors and SBFI compared to SignatureAnalyzer with L2 priors. Sensitivity is the proportion of true signatures for which there is an estimated signature with cosine similarity >>0.9. Precision is the proportion of estimated signatures for which there is a true signature with cosine similarity >>0.9.
Data are simulated as in Section C.5.1 with an expected 100 mutations per sample per signature. With N ranging from 1 to 20, sparsity ranges from 3-87% sparse, where sparsity decreases with rank (Figure C.5A). In a similar manner, the median mutational count increases with rank, ranging from 0-15 mutations (Figure C.5B). This was done to maintain a relatively constant power to discover each signature (as was done in the primary analysis with expected 1000 per sample per signature).
Figure C.5C compares rank accuracy, precision, and sensitivity between bayesNMF and SignatureAnalyzer. Our Poisson-Truncated Normal + MH with SBFI performs comparably to SignatureAnalyzer with L2 priors. However, for large ranks, SBFI might induce too much sparsity causing precision and sensitivity to drop below SignatureAnalyzer.
This sensitivity analysis shows that even with sparse data, bayesNMF is able to learn ranks between 1 and 10 with comparable precision and sensitivity to SignatureAnalyzer, though it is prone to underestimating high ranks more than SignatureAnalyzer.
C.6 Sensitivity Analysis: Range of Learned Rank
In the main text simulations learning rank, we specify the range 1-20 to bayesNMF. In this sensitivity analysis, we inspect performance when this range is expanded to 1-40 or narrowed to 1-10. We show that expanding the range does not bias learned rank upwards, and similarly that narrowing the range simply truncates estimated rank but does not bias estimated rank downward when the true rank is in the provided range.
Figure C.6: Rank bias, precision, sensitivity, and time for sensitivity analyses adjusting rank range. A. Results for rank = 1:40. B. Results for rank = 1:10.
When the range is expanded to consider ranks 1-40, we see the same general results (with a bit of noise) in terms of rank bias, precision, and sensitivity (Figure C.6A). This sensitivity analysis shows that increasing the maximum rank does not bias the learned rank upwards. It also shows that our model’s underestimation of ranks 16-20 in the primary analysis is due to a lack of signal in the data, not a downward bias from specifying a maximum rank of 20.
When the range is narrowed to consider ranks 1-10, we see the same general results (with a bit of noise) for true ranks 1-10 in terms of rank bias, precision, and sensitivity (Figure C.6B). This means that decreasing the maximum rank does not bias the learned rank downwards, as long as the true rank is within the specified range. For ranks over 10, bayesNMF estimates the specified maximum rank of 10. This shows that when there is enough signal in the data, bayesNMF is able to learn that maximum rank if appropriate, reassuring us that our model’s underestimation of ranks 16-20 is due to a lack of signal in the data, not a downward bias from the specified maximum rank. Further, this leads us to the conclusion that if the maximum rank is estimated, it is recommended to rerun the sampler with a higher maximum.
Appendix D Appendix D: Additional Data Application Results
D.1 Identifying Hypermutated Samples
The total number of mutations per sample, $m_{g}=\sum_{k}M_{kg}$, is modeled by a negative binomial mixture model, which is fit using the Expectation-Maximization algorithm [Dempster et al., 1977]. The number of mixture components, $C\in\{1,2,...,10\}$ is chosen in a two-step fashion: first, if BIC is optimized at 1 component, we say there are no hypermutated samples, and otherwise, we choose the number of components between 2 and 10 to optimize the silhouette score. Samples are assigned to clusters based on the mixture component with the highest a-posteriori probability. Figure D.1 shows the distribution of $m_{g}$ for each histology group as well as the final clustering into hypermutated and non-hypermutated samples.
Figure D.1: PCAWG data: total mutations per sample $m_{g}$ (log10 scale) plotted for each histology group and colored by hypermutation status. The y-axis labels report the number of cancer genomes $G$ separately for non-hypermutated (black) and hypermutated (red).
D.2 Additional Results
For PCAWG results using bayesNMF+SBFI and SignatureAnalyzer+ARD, Figure D.2 reports median number of mutations attributed to each signature and the cosine similarity of each estimated signature to its assigned reference signature (posterior average for bayesNMF).
Table D.5 reports estimated rank for PCAWG results using bayesNMF+SBFI and SignatureAnalyzer+ARD.
Figure D.3 investigates signatures assigned by both bayesNMF and SignatureAnalyzer versus signatures assigned by bayesNMF only. The signatures assigned by bayesNMF only have lower cosine similarity to the assigned reference and a lower proportion of assignment votes, indicating uncertainty in assignment.
Figure D.2: Results of bayesNMF+SBFI and SignatureAnalyzer+ARD on PCAWG SBS mutations data. Signatures identified in each histology group, sized by cosine similarity to the COSMIC reference signature and colored by median contribution among samples with nonzero mutations attributed to the signature (grey if median <<1). A. bayesNMF+SBFI. B. SignatureAnalyzer+ARD. (Figures inspired by Alexandrov et al. [2020] Figure 3.)
Name SignatureAnalyzer Rank bayesNMF Rank
Biliary AdenoCA 12 9
Bladder TCC 13 9
Bone Epith 5 4
Bone Osteosarc 9 7
Breast AdenoCA 15 8
CNS GBM 7 6
CNS Medullo 5 4
CNS Oligo 4 4
CNS PiloAstro 2 3
Cervix SCC 6 4
ColoRect AdenoCA 27 14
Eso AdenoCA 22 10
Head SCC 18 9
Kidney ChRCC 4 4
Kidney RCC 10 7
Liver HCC 27 11
Lung AdenoCA 22 11
Lung SCC 20 11
Lymph BNHL 17 10
Lymph CLL 7 5
Myeloid AML 3 2
Myeloid MPN 6 4
Ovary AdenoCA 14 9
Panc AdenoCA 18 9
Panc Endocrine 9 6
Prost AdenoCA 13 5
Skin Melanoma 76 10
SoftTissue Leiomyo 5 4
SoftTissue Liposarc 4 4
Stomach AdenoCA 27 11
Thy AdenoCA 3 3
Uterus AdenoCA 17 10
Table D.5: Estimated ranks of bayesNMF+SBFI and SignatureAnalyzer+ARD on PCAWG SBS mutations data.
Figure D.3: Uncertainty results on PCAWG data, separating signatures identified by both bayesNMF and SignatureAnalyzer and those identified by bayesNMF alone. A. Cosine similarity to assigned reference signature. B. Voting weight to assigned reference signature.

### v3

@@
 Abstract
-Bayesian Non-Negative Matrix Factorization (NMF) is a method of interest in fields including genomics, neuroscience, and audio and image processing.
-Bayesian Poisson NMF is of particular importance for counts data, for example in cancer mutational signature analysis.
-However, MCMC methods for Bayesian Poisson NMF require a computationally intensive augmentation.
-Further, identifying latent rank is necessary, but commonly used heuristic approaches are slow and potentially subjective, while methods that learn rank automatically are unable to provide posterior uncertainties.
-In this paper, we introduce bayesNMF, a computationally efficient Gibbs sampler for Bayesian Poisson NMF.
-Metropolis-Hastings steps are used to avoid augmentation, where full conditionals from a Normal-likelihood NMF is used as geometry-informed, high-overlap proposals.
-We additionally define sparse Bayesian factor inclusion (SBFI) as a method to identify rank automatically while providing posterior uncertainty quantification.
-We provide an open-source R software package with all of the models and plotting capabilities demonstrated in this paper on GitHub at jennalandy/bayesNMF, and supplemental materials are available online.
+Bayesian Poisson Non-Negative Matrix Factorization (NMF) is widely used to model count data, including in cancer mutational signature analysis.
+However, standard Gibbs samplers rely on computationally expensive Poisson augmentation, and current software implementations learn the latent rank either through slow and potentially subjective heuristic rank selection or with automatic approaches that do not report posterior uncertainty.
+In this paper, we introduce bayesNMF, an MH-within-Gibbs sampler to address both of these limitations.
+First, we define high-overlap proposals for Metropolis-Hastings sampling to remove the need for Poisson augmentation.
+Second, we define a BIC-based sparsity prior to learn rank automatically within the Bayesian formulation while allowing for posterior uncertainty quantification.
+We provide an open-source R software package with all of the models and plotting capabilities demonstrated in this paper on GitHub at jennalandy/bayesNMF.
 Although our applications focus on cancer mutational signatures, our software and results can be extended to any use of Bayesian Poisson NMF.
@@
 1 Introduction
-Non-Negative Matrix Factorization (NMF) is a widely used unsupervised technique for dimension reduction and intuitive parts-based representation [Lee and Seung, 1999].
-In cancer genomics, Poisson NMF can be applied to mutation counts to model mutational processes with latent factors, assuming that multiple processes contribute additively to tumor genomes [Alexandrov et al., 2013].
-Bayesian Poisson NMF allows for uncertainty quantification and the integration of prior knowledge [Cemgil, 2009, Schmidt et al., 2009].
-The computational cost of MCMC methods (e.g., Gibbs samplers) for Bayesian NMF grows with the number of samples $G$ (columns), latent factors $N$, and variables $K$ (rows).
-With a Normal likelihood, conditional updates take the form of standard distributions, requiring $(K+G)\times N$ Gibbs updates per iteration, plus updates for residual variance [Schmidt et al., 2009].
-In contrast, Bayesian NMF with a Poisson likelihood involves a data augmentation step, called Poisson augmentation, that decomposes each observed count into latent factor-specific counts, introducing $K\times G\times N$ auxiliary latent variables.
-This results in $(K+G+KG)\times N$ Gibbs updates [Cemgil, 2009], leading to a substantially higher computational cost.
-With growing sources of publicly available data for the analysis of mutational signatures [Consortium, 2010, Weinstein et al., 2013, Consortium, 2020] and a growing reference set of previously discovered signatures [Alexandrov et al., 2020], computational efficiency is a key consideration for work in this field.
-In addition, identifying latent rank is a fundamental aspect of any factor model.
-Most Bayesian NMF methods take a heuristic approach: they fit a separate model for each rank and optimize regularized metrics such as the Bayesian information criterion (BIC) [Rosales et al., 2016, Gori and Baez-Ortega, 2018] or determine the “elbow” of metrics such as mean squared error (MSE) or cosine similarity [Gori and Baez-Ortega, 2018, Islam et al., 2022].
-This is a computationally intensive process and requires post hoc, and potentially subjective, model selection.
-Existing methods that automatically learn rank as part of the Bayesian model with automatic relevance determination (ARD) do so by optimizing the posterior numerically, which means they are fast but result in point estimates without posterior uncertainties [Kasar et al., 2015].
-Another recent method, compressive NMF [Zito and Miller, 2024, CompNMF], is able to learn rank automatically while maintaining posterior uncertainty, but it does not ease the computational burden of Poisson augmentation.
-Table 1 summarizes the relative strengths of existing methods for Poisson Bayesian NMF and identifies the gap that this work addresses: we introduce a Gibbs sampler for Bayesian Poisson NMF which 1) is computationally efficient by avoiding Poisson augmentation, 2) automatically determines latent rank as a part of the Bayesian model, and 3) reports posterior uncertainties.
-We avoid the computationally intensive Poisson augmentation through Metropolis Hastings sampling with high-overlap, geometry-informed proposals derived from a Normal-likelihood Bayesian NMF.
-We show the accuracy and efficiency of this approach through simulation studies and apply it to 32 cancer types from the Pan-Cancer Analysis of Whole Genomes (PCAWG) database [Consortium, 2020].
-We also provide an open-source R software package, bayesNMF, which is available on GitHub at jennalandy/bayesNMF.
+Non-Negative Matrix Factorization (NMF) is a widely used unsupervised technique for dimension reduction and intuitive parts-based representation [Lee and Seung, 1999], and is typically formulated with either a Poisson or a Normal likelihood.
+In cancer genomics, Poisson NMF applied to mutation counts assumes that tumor genomes arise from multiple additive mutational processes, with the resulting latent factors termed mutational signatures [Alexandrov et al., 2013].
+Within this framework, Bayesian NMF allows for integration of prior knowledge and posterior uncertainty quantification [Cemgil, 2009, Schmidt et al., 2009].
+The computational cost of a Gibbs sampler for Bayesian NMF grows with the number of samples $G$ (columns), latent factors $N$, and variables $K$ (rows).
+A Normal likelihood requires approximately $N(K+G)$ updates per iteration [Schmidt et al., 2009].
+A Poisson likelihood, however, relies on Poisson augmentation for conditional updates to follow standard distributions.
+Each observed count is decomposed into latent factor-specific counts, introducing $NKG$ auxiliary parameters.
+This requires $N(K+G+KG)$ updates per iteration [Cemgil, 2009], resulting in a substantially higher computational cost.
+With growing publicly available data sources [Consortium, 2010, Weinstein et al., 2013, Consortium, 2020] and an expanding set of previously discovered signatures [Alexandrov et al., 2020], computational efficiency is a central constraint in large-scale mutational signature analysis.
+Further, identifying latent rank is a fundamental aspect of any factor model.
+Most Bayesian NMF methods take a heuristic approach, fitting a separate model for each rank and optimizing regularized metrics such as the Bayesian information criterion (BIC) [Rosales et al., 2016, Gori and Baez-Ortega, 2018] or determining the “elbow” of metrics such as MSE or cosine similarity [Gori and Baez-Ortega, 2018, Islam et al., 2022].
+This process is computationally intensive and relies on post hoc and potentially subjective model selection.
+Existing methods address subsets of these challenges, but trade off posterior uncertainty, computational efficiency, and rank learning in different ways (Table 1).
+SignatureAnalyzer [Kasar et al., 2015] uses automatic relevance determination (ARD) to learn rank as part of the Bayesian model.
+It avoids Poisson augmentation by optimizing the posterior numerically, but as a result, it provides only point estimates without posterior uncertainties.
+On the other hand, SigFit [Gori and Baez-Ortega, 2018] avoids Poisson augmentation by performing posterior inference in Stan [Carpenter et al., 2017], which implements Hamiltonian Monte Carlo [Neal and others, 2011], but selects rank heuristically across candidate models.
+Compressive NMF [Zito and Miller, 2024] learns rank automatically while maintaining posterior uncertainty, but relies on computationally expensive Poisson augmentation.
+To our knowledge, no method simultaneously 1) avoids Poisson augmentation, 2) learns rank within the Bayesian formulation, and 3) reports posterior uncertainty.
+This paper introduces two innovations to address this gap.
+First, we define high-overlap proposals for Metropolis-Hastings-within-Gibbs sampling to avoid Poisson augmentation.
+Second, we define a BIC-based sparsity prior to learn rank within the Bayesian model.
+The accuracy and efficiency of the combined approach are evaluated through simulation studies, and results are presented on 32 cancer types from the Pan-Cancer Analysis of Whole Genomes (PCAWG) database [Consortium, 2020].
+Methods are implemented in an open-source R package, bayesNMF, available on GitHub at jennalandy/bayesNMF.
 2 Background
@@
 SigneR (2016) $\checkmark$
-SigFit (2018) $\checkmark$
+SigFit (2018) $\checkmark$ $\checkmark$
 CompNMF (2024) $\checkmark$ $\checkmark$
-Table 1: Software for Bayesian NMF applied to de-novo mutational signature analysis: SignatureAnalyzer [Kasar et al., 2015, Kim et al., 2016], SigneR [Rosales et al., 2016], SigFit [Gori and Baez-Ortega, 2018], and CompNMF [Zito and Miller, 2024].
-An ideal software would learn rank automatically, report posterior uncertainty, and avoid computationally intensive Poisson augmentation.
+Table 1: Software for Bayesian Poisson NMF applied to de-novo mutational signature analysis: SignatureAnalyzer [Kasar et al., 2015, Kim et al., 2016], SigneR [Rosales et al., 2016], SigFit [Gori and Baez-Ortega, 2018], and CompNMF [Zito and Miller, 2024].
+An ideal software would automatically learn rank within the Bayesian formulation, report posterior uncertainty, and avoid computationally intensive Poisson augmentation.
 2.1 Non-Negative Matrix Factorization (NMF) and Bayesian NMF
-NMF is an unsupervised method to decompose non-negative data $M$ ($K$ variables by $G$ samples) into two lower-rank non-negative matrices, $P$ and $E$ [Paatero and Tapper, 1994, Lee and Seung, 1999].
-NMF represents each observation of the data matrix, or column of $M$, as a linear combination of latent factors, or columns of $P$, with weights in $E$:
-$M_{kg}$ $=\sum_{n=1}^{N}P_{kn}E_{ng},\hskip 11.38109pt\mbox{i.e}\;\;M=PE,\hskip 11.38109ptM\in\mathbb{R}_{\geq 0}^{K\times G},P\in\mathbb{R}_{\geq 0}^{K\times N},E\in\mathbb{R}_{\geq 0}^{N\times G}.$
-NMF is typically fit by minimizing a reconstruction error, often the generalized Kullback-Leibler (KL) divergence between $M$ and $PE$ (Equation 1) or the Frobenius norm of MPEM-PE (Equation 2).
+NMF is an unsupervised method to decompose non-negative data $M$ ($K$ variables as rows by $G$ samples as columns) into two lower-rank non-negative matrices, $P$ and $E$ [Paatero and Tapper, 1994, Lee and Seung, 1999].
+NMF represents each observation of the data matrix (column of $M$) as a linear combination of latent factors (columns of $P$):
+$M_{kg}$ $=\sum_{n=1}^{N}P_{kn}E_{ng},\hskip 11.38109pt\mbox{i.e.,}\;\;M=PE,\hskip 11.38109ptM\in\mathbb{R}_{\geq 0}^{K\times G},P\in\mathbb{R}_{\geq 0}^{K\times N},E\in\mathbb{R}_{\geq 0}^{N\times G}.$
+NMF is typically fit by minimizing a reconstruction error, often the generalized Kullback-Leibler (KL) divergence (Equation 1) or the Frobenius norm (Equation 2).
+$KL(M||PE)$ $=\sum_{kg}\left(M_{kg}\log\frac{M_{kg}}{(PE)_{kg}}-M_{kg}+(PE)_{kg}\right)$ (1)
+[…]
 For each, Lee and Seung [2000] adapt gradient descent by deriving a step size that preserves nonnegativity, resulting in multiplicative update algorithms.
-$KL(M||PE)$ $=\sum_{kg}\left(M_{kg}log\frac{M_{kg}}{(PE)_{kg}}-M_{kg}+(PE)_{kg}\right)$ (1)
-[…]
-Minimizing KL-divergence is equivalent to maximizing a Poisson likelihood on $M$ centered at $PE$ [Cemgil, 2009].
-Similarly, minimizing the Frobenius norm is equivalent to maximizing a Normal likelihood on $M$ centered at $PE$ [Schmidt et al., 2009].
-Bayesian approaches add priors on $P$ and $E$ [Cemgil, 2009, Schmidt et al., 2009].
-To sample Bayesian Poisson NMF posterior distributions with standard MCMC methods, Poisson augmentation is necessary to achieve easily-sampled full conditional distributions of $P$ and $E$ [Cemgil, 2009].
-This introduces latent variables $Z_{kng}$, defined as the part of the $M_{kg}$ count attributed to latent factor $n$.
-Each $Z_{kng}$ follows a Poisson distribution centered at $P_{kn}E_{ng}$ and satisfies $M_{kg}=\sum_{n}Z_{kng}$.
-This imposes $K\cdot N\cdot G$ additional Gibbs updates at each iteration, increasing the computational requirements of Bayesian Poisson NMF far beyond that of Bayesian Normal NMF.
+These equate to probabilistic models: minimizing KL-divergence is equivalent to maximizing a Poisson likelihood on $M$ with mean $PE$ [Cemgil, 2009], and minimizing the Frobenius norm is equivalent to maximizing a Normal likelihood on $M$ with mean $PE$ [Schmidt et al., 2009].
+Bayesian approaches place priors on $P$ and $E$ [Cemgil, 2009, Schmidt et al., 2009].
+Sampling the posterior distribution of Bayesian Poisson NMF with a standard Gibbs sampler requires Poisson augmentation [Cemgil, 2009].
+This introduces latent variables $Z_{kng}$, defined as the counts from $M_{kg}$ attributed to latent factor $n$.
+Each $Z_{kng}$ follows a Poisson distribution with mean $P_{kn}E_{ng}$ and satisfies $M_{kg}=\sum_{n}Z_{kng}$.
+These additional $KNG$ parameters increase the computational requirements of Bayesian Poisson NMF far beyond those of Bayesian Normal NMF.
 2.2 Cancer Mutational Signatures
 Cancer is driven by genetic changes, including somatic DNA mutations, which can be caused by multiple processes.
-Understanding which mutational process activities are reflected in a tumor genome can improve our understanding of cancer development and answer questions about subtyping, prognosis, and treatment [Nik-Zainal et al., 2012a, b, Alexandrov et al., 2013, Rosales et al., 2016].
+Understanding which mutational process activities are reflected in a tumor genome can help answer questions about cancer development, subtyping, prognosis, and treatment [Nik-Zainal et al., 2012a, b, Alexandrov et al., 2013, Rosales et al., 2016].
 A computationally-derived mutational signature models the specific patterns produced by a mutational process.
-Examples include DNA damage repair deficiencies [Zámborszky et al., 2017], deamination of 5-methylcytosine [Nik-Zainal et al., 2012a], tobacco smoke [Alexandrov et al., 2013, Nik-Zainal et al., 2015], and ultraviolet radiation [Nik-Zainal et al., 2015, Saini et al., 2016, Hayward et al., 2017].
+Signatures have previously been linked to DNA damage repair deficiencies [Zámborszky et al., 2017], deamination of 5-methylcytosine [Nik-Zainal et al., 2012a], tobacco smoke [Alexandrov et al., 2013, Nik-Zainal et al., 2015], and ultraviolet radiation [Nik-Zainal et al., 2015, Saini et al., 2016, Hayward et al., 2017].
 The Catalog of Somatic Mutations in Cancer (COSMIC) [Tate et al., 2019] provides a high-confidence set of reference signatures.
@@
 […]
-Mutational signature analysis is a rapidly growing field driven by an increasing number of publicly available databases, including COSMIC [Tate et al., 2019], mSignatureDB [Huang et al., 2018], the Pan-Cancer Analysis of Whole Genomes (PCAWG) [Consortium, 2020], and the compendium of Mutational Signatures of Environmental Agents [Kucab et al., 2019].
+Mutational signature analysis is a rapidly growing field driven by an increasing number of publicly available databases, including COSMIC [Tate et al., 2019], mSignatureDB [Huang et al., 2018], PCAWG [Consortium, 2020], and the compendium of Mutational Signatures of Environmental Agents [Kucab et al., 2019].
 With increased data sizes and availability, computational efficiency must be carefully considered when making modeling and implementation choices.
@@
 3.1 MH-within-Gibbs for Computational Efficiency
-We introduce two efficient models for Poisson Bayesian NMF. The first uses a Truncated Normal (TN) prior on each element of $P$ and $E$:
+We introduce two efficient models for Poisson Bayesian NMF. The first uses a Truncated Normal prior on each element of $P$ and $E$:
 Likelihood: $\quad M_{kg}\sim\text{Poisson}((PE)_{kg})$ (3)
-Priors: $\quad P_{kn}\sim\text{TN}(\mu^{P}_{kn},\sigma^{P}_{kn},0,\infty),\quad E_{ng}\sim\text{TN}(\mu^{E}_{ng},\sigma^{E}_{ng},0,\infty).$ (4)
+Priors: $\quad P_{kn}\sim\text{TruncNorm}(\mu^{P}_{kn},\sigma^{P}_{kn},0,\infty),\quad E_{ng}\sim\text{TruncNorm}(\mu^{E}_{ng},\sigma^{E}_{ng},0,\infty).$ (4)
 Specifications of hyperprior distributions and hyperparameters can be found in Appendix A.
@@
 Each term $(P^{*}E)_{kg}$ depends on $P^{*}_{kn}$ for all $n$, linking likelihoods across $n$, and preventing factorization into a standard family:
-$f(P_{kn}^{*}|P_{-kn},E,M)$ $\propto p_{\text{TN}}(P_{kn}^{*})\prod_{g}p_{\text{Poisson}}\left(M_{kg}|\lambda=(P^{*}E)_{kg}\right)$ (5)
-$\propto p_{\text{TN}}(P^{*}_{kn})e^{-\sum_{g}(P^{*}E)_{kg}}\prod_{g}(P^{*}E)_{kg}^{M_{kg}}.$
+$f(P_{kn}^{*}|P_{-kn},E,M)$ $\propto p_{\text{TruncNorm}}(P_{kn}^{*})\prod_{g}p_{\text{Poisson}}\left(M_{kg}|\lambda=(P^{*}E)_{kg}\right)$ (5)
+$\propto p_{\text{TruncNorm}}(P^{*}_{kn})e^{-\sum_{g}(P^{*}E)_{kg}}\prod_{g}(P^{*}E)_{kg}^{M_{kg}}.$
 Poisson augmentation resolves this in previous methods [Cemgil, 2009], but at high computational cost.
@@
 We construct proposal distributions $g$ from the full conditional form of a Normal Bayesian NMF model with identical Truncated Normal priors to our target model:
-$g(P^{*}_{kn}$ $|P_{-kn},E,M)\propto p_{\text{TN}}(P^{*}_{kn})\prod_{g}p_{\text{Normal}}\left(M_{kg}|\mu=(P^{*}E)_{kg},\sigma^{2}=(PE)_{kg}\right)$ (6)
-$\sim\text{Truncated Normal}\left(P^{*}_{kn}|\mu=m(M,P,E,\theta),\sigma=s(M,P,E,\theta),0,\infty\right)$
+$g(P^{*}_{kn}|P_{-kn},E,M)$ $\propto p_{\text{TruncNorm}}(P^{*}_{kn}|\theta)\prod_{g}p_{\text{Normal}}\left(M_{kg}|\mu=(P^{*}E)_{kg},\sigma^{2}=(PE)_{kg}\right)$ (6)
+$=p_{\text{TruncNorm}}\left(P^{*}_{kn}|\mu=m(M,P,E,\theta),\sigma=s(M,P,E,\theta),0,\infty\right)$
 where $\theta$ holds prior and hyperprior parameters. The Normal model is never sampled directly. Specifications of the location $m(\cdot)$ and scale $s(\cdot)$ functions are in Appendix A.
-Instead of modeling variances $\sigma^{2}_{kg}$ as would be required in a Normal Bayesian NMF, we substitute it with $(PE)_{kg}$ in the proposal.
+Instead of modeling variances $\sigma^{2}_{kg}$ as would be required in a Normal Bayesian NMF, we substitute $(PE)_{kg}$ into the proposal.
 This allows the proposal to reflect the Poisson mean-variance relationship of the target model without additional parameters.
-This only affects the proposal $g$, which needs to be a proper distribution but does not need to be an actual full conditional to maintain detailed balance.
+This affects only the proposal $g$, which needs to be a proper distribution in order to maintain detailed balance, but does not need to be an actual full conditional.
 After drawing a proposed value $P^{*}_{kn}$, the acceptance ratio can be computed as a ratio of likelihoods because priors cancel out from both the target and proposal distributions:
@@
 $=\prod_{g}\frac{p_{\text{Poisson}}\left(M_{kg}|\lambda=(P^{*}E)_{kg}\right)\cdot p_{\text{Normal}}\left(M_{kg}|\mu=(PE)_{kg},\sigma^{2}=(P^{*}E)_{kg}\right)}{p_{\text{Poisson}}\left(M_{kg}|\lambda=(PE)_{kg}\right)\cdot p_{\text{Normal}}\left(M_{kg}|\mu=(P^{*}E)_{kg},\sigma^{2}=(PE)_{kg}\right)}.$ (8)
-For a fixed rank, the Normal and Poisson Bayesian NMF models have identical MAP solutions (up to scaling and permutation symmetries of NMF [Lee and Seung, 1999, Donoho and Stodden, 2003]).
-Consequently, the Normal-based proposal captures the local curvature of the Poisson posterior near its mode, ensuring high overlap.
-This proposal is analogous to gradient-based methods (e.g., Metropolis-adjusted Langevin [Roberts and Tweedie, 1996] or Hamiltonian Monte Carlo [Neal et al., 2011]), but we do not compute gradients explicitly.
-The second model we introduce uses an Exponential prior.
-The proposal still takes the form of a Truncated Normal distribution, though functions $m(\cdot)$ and $s(\cdot)$ change.
+For a fixed rank, the Normal and Poisson Bayesian NMF models have approximately the same MAP solutions (up to scaling and permutation symmetries of NMF [Lee and Seung, 1999, Donoho and Stodden, 2003]).
+Consequently, the Normal-based proposal captures the local curvature of the Poisson posterior near its mode, designed to have high overlap.
+The second model we introduce places an Exponential prior on elements of $P$ and $E$.
+In order to adapt the Poisson+MH approach to this model, the proposal distribution is now based on a paired Normal likelihood NMF with Exponential priors.
+This proposal again takes the form of a Truncated Normal distribution but with redefined functions $m(\cdot)$ and $s(\cdot)$.
 In general, the Poisson+MH approach can only be used when the prior (which we assume is shared by the target and the proposal) leads to Normal Bayesian NMF full conditionals that are of known families, meaning Gamma priors cannot be used with this setup.
-The limitation is practical rather than fundamental—other proposal families could in principle accommodate Gamma priors but are not considered here.
+The limitation is practical rather than fundamental—other proposal families could, in principle, accommodate Gamma priors but are not considered here.
 3.2 Sparse Bayesian Factor Inclusion (SBFI) for Learning Rank
 Previous Bayesian work by Grabski et al.
-[2023a, b] in the multi-study setting learns factor sharing patterns and total latent dimension using a binary factor inclusion matrix $A$.
+[2023, 2025] in the multi-study setting learns factor sharing patterns and total latent dimension using a binary factor inclusion matrix $A$.
 We adapt this approach to learn rank automatically in our single-study setting and introduce a new prior and hyperprior structure such that the induced prior on the expected rank is approximately uniform.
-We term this approach Bayesian factor inclusion (BFI).
+We refer to this as Bayesian factor inclusion (BFI).
 BFI uses a diagonal binary factor inclusion matrix, $A$, where $A_{nn}=1$ if latent factor $n$ is included in the model, and 0 if excluded, redefining NMF as
-$\mathbb{E}[M_{kg}]=\sum_{n=1}^{N}P_{kn}A_{nn}E_{ng},\quad\quad\mbox{i.e.}\quad\quad\mathbb{E}[M]=PAE$
-with latent rank $N^{\prime}=\sum_{n=1}^{N}A_{nn}$. When $A_{nn}=0$, hyperpriors of $P$ and $E$ retain some information on the excluded signature, reducing the extent of label switching.
+$\mathbb{E}[M_{kg}]=\sum_{n=1}^{N}P_{kn}A_{nn}E_{ng},\quad\quad\mbox{i.e.,}\quad\quad\mathbb{E}[M]=PAE$
+with a prespecified large $N$ and learned latent rank $N^{\prime}=\sum_{n=1}^{N}A_{nn}$.
+If $A$ is fixed as an identity matrix, this formulation is identical to that in Section 3.1 for latent rank $N$.
 Each factor inclusion indicator $A_{nn}$ has a Bernoulli prior.
-The prior probability of factor inclusion $q$ is determined by a random variable $R$, which represents the expected rank and has a uniform hyperprior between 0 and $N$.
+The prior probability of factor inclusion $q$ is determined by a random variable $R$ representing the expected rank with a uniform hyperprior.
 The value $q$ is bounded away from 0 and 1 to prevent the Gibbs sampler from getting stuck at the endpoint ranks 0 or $N$:
@@
 $p(R=r)=1/(N+1)\text{ for }r=0,...,N.$
-In practice, we set $N$ to a large number. By learning the posterior distribution of $A$, we also learn the implied posterior distribution of the latent rank $N^{\prime}=\sum_{n=1}^{N}A_{nn}$.
+If $A_{nn}$ is sampled as 0, hyperpriors of $P$ and $E$ retain some information on the excluded signature, reducing the extent of label switching.
 Including $A$ in the model induces a highly multimodal posterior.
 We adapt the tempering approach used in Grabski et al.
-[2023b] for the Gibbs updates of $A$ and $R$, where the likelihood term is raised to a temperature power $\gamma$: $A_{nn}\sim p(A_{nn})p(M|A_{nn},...)^{\gamma}$ and $R\sim p(R)p(M|R,...)^{\gamma}$.
+[2025] for the Gibbs updates of $A$ and $R$, where the likelihood term is raised to a temperature power $\gamma$: $A_{nn}\sim p(A_{nn})p(M|A_{nn},...)^{\gamma}$ and $R\sim p(R)p(M|R,...)^{\gamma}$.
 The temperature $\gamma$ progresses from 0 to 1 over the course of a burn-in period.
@@
 No tempered samples are used for posterior inference.
-We also introduce Sparse Bayesian factor inclusion (SBFI), which replaces the likelihood in the Gibbs update of $A_{nn}$ with $e^{-\frac{1}{2}\text{BIC}}$ to encourage sparsity in the latent rank:
-BIC $=-2\cdot\text{log}\left[p_{\text{Poisson}}(M|A_{nn}=a,...)\right]+\text{log}(G)\cdot(K+G)N^{(a)}$
-[…]
-where $N^{(a)}=\sum_{n^{\prime}}A_{n^{\prime}n^{\prime}}$ when $A_{nn}=a$, $G$ is the sample size, and $(K+G)N^{(a)}$ is the number of functional parameters in $P$ and $E$.
-Expanding $e^{-\frac{1}{2}\text{BIC}}$ shows that, once $\gamma=1$, SBFI is equivalent to a Gibbs update under the following sparse prior on $A_{nn}$:
-$p_{\text{SBFI}}(A_{nn}=1)$ $=\frac{qG^{-\frac{1}{2}(K+G)}}{qG^{-\frac{1}{2}(K+G)}+(1-q)}.$
+We also introduce Sparse BFI (SBFI) with a sparse prior on $A_{nn}$:
+$p$ $_{\text{SBFI}}(A_{nn}=1)=\frac{qG^{-\frac{1}{2}(K+G)}}{qG^{-\frac{1}{2}(K+G)}+(1-q)}$
+$p$ $_{\text{SBFI}}(A_{nn}=a|...)\propto q^{a}(1-q)^{1-a}\cdot\left[p(M|A_{nn}=a,...)\cdot G^{-\frac{1}{2}(K+G)N^{(a)}}\right]^{\gamma}$ (9)
+where $G$ is sample size, $N^{(a)}=\sum_{n^{\prime}}A_{n^{\prime}n^{\prime}}$ when $A_{nn}=a$, and $(K+G)N^{(a)}$ is the number of functional parameters in $P$ and $E$.
+During the tempering phase of the sampler, tempering is also applied to the sparsity-inducing term to avoid early convergence on a low rank before signatures are learned.
+Again, no tempered samples are used for posterior inference.
+The structure of this prior results in a BIC-like penalty on rank.
+Specifically, once tempering is complete, the Gibbs update in Equation 3.2 is identical to the product of the standard BFI prior and the term $e^{-\frac{1}{2}\text{BIC}}$:
+$\text{BIC}=-2\cdot\text{log}\left[p_{\text{Poisson}}(M|A_{nn}=a,...)\right]+\text{log}(G)\cdot(K+G)N^{(a)},$
+$q^{a}(1-q)^{1-a}\cdot e^{-\frac{1}{2}\text{BIC}}=q^{a}(1-q)^{1-a}\cdot p(M|A_{nn}=a,...)\cdot G^{-\frac{1}{2}(K+G)N^{(a)}}.$
 3.3 Baseline Models
@@
 The Poisson-Gamma model is available in SigneR [Rosales et al., 2016] while Truncated Normal and Exponential priors as well as Normal likelihoods are available in SignatureAnalyzer [Kasar et al., 2015, Kim et al., 2016].
-We implement our own samplers to enable direct models comparisons where all settings but likelihoods and priors are held constant.
+We implement our own samplers to enable direct model comparisons where all settings but likelihoods and priors are held constant.
 We do not implement a standard Gibbs sampler for the Poisson-Truncated Normal model because the full conditionals are not from distributions that can be directly sampled, even with Poisson augmentation.
 The Poisson-Gamma model is the closest comparison.
-To evaluate the ability of SBFI to learn rank, we compare it to minBIC, a procedure which fits a separate Poisson+MH sampler for each candidate rank and selects the one minimizing BIC.
+To evaluate the ability of SBFI to learn rank, we compare it to minBIC, a procedure that fits a separate Poisson+MH sampler for each candidate rank and selects the one minimizing BIC.
 We also compare it to automatic relevance determination as implemented in SignatureAnalyzer, which learns rank but does not report posterior uncertainty.
-Likelihood Prior Sampler Purpose Reference
-Poisson Truncated Normal MH-within-Gibbs This Work -
-Poisson Exponential MH-within-Gibbs This Work -
-Poisson Gamma Standard Gibbs Baseline SigneR
-Poisson Exponential Standard Gibbs Baseline SignatureAnalyzer
-Normal Truncated Normal Standard Gibbs Motivation SignatureAnalyzer
-Normal Exponential Standard Gibbs Motivation SignatureAnalyzer
+Under the non-negativity constraints in NMF, SignatureAnalyzer’s L1 and L2 priors correspond to Exponential and Truncated Normal priors, respectively.
+Likelihood Prior Sampler Purpose
+Poisson Truncated Normal MH-within-Gibbs This Work
+Poisson Exponential MH-within-Gibbs This Work
+Poisson Gamma Standard Gibbs Baseline
+Poisson Exponential Standard Gibbs Baseline
+Normal Truncated Normal Standard Gibbs Motivation
+Normal Exponential Standard Gibbs Motivation
 Table 2: List of models implemented in bayesNMF.
@@
 Standard Normal samplers serve as motivation for the proposal distributions of MH-within-Gibbs models.
-References refer to other software implementing the same likelihood-prior combinations, though they are re-implemented in bayesNMF to allow for direct comparisons.
-(SignatureAnalyzer [Kasar et al., 2015, Kim et al., 2016], and SigneR [Rosales et al., 2016].)
 3.4 Inference
-Inference is performed on the final 1000 posterior samples.
-When learning rank, we first compute $\hat{A}$, the maximum a-posteriori (MAP) estimate of the signature inclusion matrix, at its mode.
-We then subset inference samples to those that match this MAP.
-To summarize the inference samples, we next remove the scale indeterminacy of NMF by normalizing each iteration’s $P^{(i)}$ and $E^{(i)}$ matrices for columns of $P^{(i)}$ to sum to one.
-Then, final estimates $\hat{P}$ and $\hat{E}$ and 95% credible intervals are element-wise means and 2.5th and 97.5th percentiles, respectively.
-The estimated reconstruction $\hat{M}$ is the product $\hat{P}\hat{A}\hat{E}$.
-To compare results with a reference $P$, we assign each inference samples’s signature matrix $P^{(i)}$ to $P$ with the Hungarian algorithm [Kuhn, 1955].
+For results in Sections 5 and 6, inference is performed on the final 1000 posterior samples, though our software allows this number to be user-specified and for all samples to be stored.
+To summarize the posterior distribution when rank is fixed, we first remove the scale indeterminacy of NMF by normalizing each iteration’s $P^{(i)}$ and $E^{(i)}$ matrices so that columns of $P^{(i)}$ sum to one.
+Then, posterior means $\hat{P}$ and $\hat{E}$ are element-wise averages, and 95% credible intervals are element-wise 2.5th and 97.5th percentiles.
+Parameters $P$ and $E$ are only interpretable in the context of the inclusion structure $A$.
+Therefore, when learning rank, the target for inference is the posterior distribution conditional on the maximum a-posteriori (MAP) inclusion matrix $\hat{A}$, identified as the mode across samples $A^{(i)}$.
+Summaries of the conditional posterior are then computed from samples where $A^{(i)}=\hat{A}$.
+Within the selected samples, $\hat{P}$ and $\hat{E}$ are computed as above.
+For optional post-hoc comparison with a reference $P$, we assign each $P^{(i)}$ to $P$ with the Hungarian algorithm [Kuhn, 1955].
 This assumes each reference signature may only be assigned to a single estimated signature, and vice versa.
-We ensemble assignments over inference samples with majority voting (details in Appendix B.4).
+We ensemble assignments across samples with majority voting (details in Appendix B.4).
 4 Implementation and R Software Package
-4.1 Implementation
-The bayesNMF R software package implements an automated convergence detection algorithm and reports convergence metrics in a log file and with trace plots (Figure 1A).
-This makes the software simple to use, reduces subjectivity of sampler stopping, and ensures all models reach a similar degree of convergence so results can be compared fairly.
-Full details of model specifications, Gibbs updates, and hyperparameters are available in Appendix A.
-Details of convergence, inference, and reference assignment are in Appendix B.
+4.1 Software
+The bayesNMF R software package is available for download on GitHub at jennalandy/bayesNMF.
+Details of model specifications, Gibbs updates, and hyperparameters are available in Appendix A.
+Details of automated convergence detection and diagnostics, inference, and reference assignment are in Appendix B.
 4.2 Modeling Capabilities
-The bayesNMF R software package allows users to fit all models described in this paper with the bayesNMF function.
+Users are able to fit all models described in this paper with the bayesNMF function.
 Model specifications can be adjusted by the likelihood, prior, and MH parameters.
@@
 Figure 2: Relative performance of Bayesian NMF models.
-Grey fill represents Gamma prior on standard Poisson models, but Truncated Normal prior on Poisson+MH or Normal models.
+Grey fill represents Gamma priors on standard Poisson models, but Truncated Normal priors on Poisson+MH or Normal models.
 White fill always indicates Exponential priors.
@@
 Agreement of $\hat{E}$ (left) and $\hat{P}$ (right) between models as signature-wise minimum cosine similarities.
-Top: standard Poisson vs Normal; Middle: standard Poisson vs Poisson+MH; Bottom: standard Poisson vs standard Poisson (null case).
 Excluding 0 to 7 values per density for similarity <<0.95 for visual clarity (see Appendix Table C.1).
@@
 Performance of novel Poisson+MH models.
+Metrics for all models are available in Appendix Figure C.1.
 5 Simulation Studies
 5.1 Metrics
-Before computing metrics, we align signatures as described in Section 3.4. SignatureAnalyzer does not report posterior samples so results are aligned on final estimates only.
-When rank is known and two models are compared to one another, we report the cosine similarity of the worst-aligned signature, as well as that of the worst-aligned exposure profile.
+Before computing metrics, signatures are aligned to ground truth.
+Results for bayesNMF are aligned as described in Section 3.4.
+SignatureAnalyzer does not report posterior samples, so its results are aligned using point estimates only.
+When rank is known, and two models are compared to one another, we report the cosine similarity of the worst-aligned signature, as well as that of the worst-aligned exposure profile.
 If the two solutions are identical, both values will be exactly 1.
 We also report relative efficiency as the ratio of total runtime between two models.
-When rank is known and a model is compared to ground truth, we first evaluate NMF solutions in terms of their ability to reconstruct the original data matrix using RMSE and KL-Divergence between $M$ and $\hat{M}$.
+When rank is known, and a model is compared to ground truth, we first evaluate NMF solutions in terms of their ability to reconstruct the original data matrix using RMSE and KL-Divergence between $M$ and $\hat{M}=\hat{P}\hat{E}$.
 We then measure their ability to reconstruct ground truth latent factors using the cosine similarity of the worst-reconstructed signature.
-Finally, we report total runtime in minutes to assess computational efficiency.
-When learning rank, we evaluate NMF solutions by the bias of estimated rank as well as post-alignment sensitivity and precision.
+Finally, we report the total runtime in minutes to assess computational efficiency.
+When learning rank, we evaluate NMF solutions by the bias of the estimated rank as well as post-alignment sensitivity and precision.
 Sensitivity is the proportion of true signatures that have an estimated match of cosine similarity >>0.9.
@@
 5.2 Poisson+MH efficiently samples the correct posterior
-In this section, we further justify our Normal model-based proposals, and we show that Poisson+MH models more efficiently sample the same posterior as standard Poisson models.
-We consider true latent ranks $N\in[2,4,8,16]$ and sample sizes $G\in[16,32,64,128]$ such that $G\geq 8N$.
-For each combination, 10 datasets are generated.
-$N$ signatures are sampled from COSMIC SBS v3.3.1 [Tate et al., 2019], each normalized to sum to 1.
-The expected number of mutations for a sample $g$ is $m_{g}\sim\text{Negative Binomial}(r\cdot N,p)$}, with $r=111.11$ and $p=0.1$.
-This yields $E[m_{g}]=1000\cdot N$, or $1000$ mutations per signature per sample, on average, so that each simulation provides comparable relative power to learn each signature.
-Columns of the exposures matrix, $E$, are generated $E_{g}\sim\text{Multinomial}(m_{g},\mathbf{p})$, where $\mathbf{p}\sim\text{Dirichlet}(\mathbf{1}_{N})$.
-Mutational counts are simulated as $M_{kg}\sim\text{Poisson}(\lambda_{kg}=(PE)_{kg})$.
-Figure 2A compares point estimates.
-In all comparisons, agreement is >>0.99 in at least 90% of cases for $\hat{E}$ and 84% for $\hat{P}$.
-The top row shows Normal Bayesian NMF results in point estimates nearly identical to Poisson Bayesian NMF, justifying our Normal model-based MH proposals.
-The middle row shows the addition of MH steps does not alter point estimates of Poisson Bayesian NMF.
-The bottom row is a null case, comparing two standard Poisson samplers, to show that some variation is expected due to the randomness of Gibbs samplers.
+In this section, we treat rank as known and perform model comparisons to further justify our Normal model-based proposals and to show that Poisson+MH models more efficiently sample the same posterior as standard Poisson models.
+We consider true latent ranks $N\in[2,4,8,16]$ and sample sizes $G\in[16,32,64,128]$.
+For each $N,G$ where $G\geq 8N$, 10 datasets are generated.
+Signatures are sampled from COSMIC SBS v3.3.1 [Tate et al., 2019], each normalized to sum to 1.
+The expected total mutation count for sample $g$ is $m_{g}\sim\text{Negative Binomial}(111.11N,0.1)$}, yielding $E[m_{g}]=1000N$, or $1000$ mutations per signature per sample.
+This gives each simulation comparable relative power to learn each signature, and this explains why reconstruction errors increase with $N$.
+Columns of the exposures matrix, $E$, are generated $E_{g}\sim\text{Multinomial}(m_{g},\mathbf{p}_{g})$, where $\mathbf{p}_{g}\sim\text{Dirichlet}(\mathbf{1}_{N})$.
+Mutation counts are simulated as $M_{kg}\sim\text{Poisson}((PE)_{kg})$.
+Normal Bayesian NMF results in point estimates nearly identical to Poisson Bayesian NMF, justifying our Normal model-based MH proposals (Figure 2A, top row).
+Further, the addition of MH steps does not alter point estimates of Poisson Bayesian NMF (middle row).
+These comparisons yield agreement >>0.99 in at least 84% of cases and are very close to the null case comparing two independent chains of the standard Poisson sampler (bottom row).
 Along with aligned point estimates, the addition of MH steps increases effective sample sizes (Appendix C.1) and does not substantially change credible interval widths (Appendix C.3).
-Together, these results show the correctness of our novel Poisson+MH samplers in that they explore the same posterior space as the standard Poisson samplers.
-Figure 2B reports relative efficiency of Poisson+MH compared to standard Poisson samplers. Poisson+MH models are 3-30x faster, and the gap grows with dimensionality.
-Figure 2C reports metrics of the Poisson+MH samplers as a baseline for the comparisons of Figures 2A-B. Metrics for all models are available in Appendix Figure C.1.
+Together, these results show that our novel Poisson+MH samplers yield the same posterior inference as the standard Poisson samplers.
+Additionally, Poisson+MH models are 3-30x faster, and this gap grows with dimensionality (Figure 2B).
 5.3 SBFI learns rank
-In this section, we demonstrate the utility of SBFI for automatically learning rank.
-SBFI performs comparably to, or better than, SignatureAnalyzer’s automatic relevance determination (ARD) and reports posterior uncertainty.
-Our recommendation is bayesNMF with SBFI for Truncated Normal priors or bayesNMF with minBIC for Exponential priors.
+In this section, we demonstrate that SBFI performs comparably to or better than SignatureAnalyzer’s automatic relevance determination (ARD) while reporting posterior uncertainty.
 Figure 3: Rank bias, precision, sensitivity, and time (log scale) of rank learning approaches: bayesNMF with SBFI (upwards triangle) and minBIC (square), as well as SignatureAnalyzer with ARD (downwards triangle), each with Truncated Normal (filled) and Exponential (empty) priors.
@@
 Precision is the proportion of estimated signatures for which there is a true signature with cosine similarity >>0.9.
-Data are simulated as in Section 5.2, but with $G=64$ and true ranks $N$ between $1$ and $20$.
-10 datasets are simulated per $N$ such that no two included signatures have cosine similarity $\geq$0.8.
-We note that lower precision and sensitivity at the larger ranks may reflect the difficulty of the problem rather than flaws of the methods considered.
+Data are simulated as in Section 5.2 with $G=64$ and 10 datasets for each true rank $N$ between $1$ and $20$.
+For each dataset, no two signatures have a cosine similarity $\geq$0.8.
+Reduced precision and sensitivity observed at larger ranks likely reflect the difficulty of the problem (i.e., large $N$ relative to $G$), rather than deficiencies of the methods considered.
 Figure 3 reports rank bias, precision, sensitivity, and runtime for each model as true rank increases along the x-axis.
-For a Truncated Normal / L2 prior (filled shapes), all approaches estimate rank very well, but with different error patterns.
-SignatureAnalyzer overestimates low and high ranks, while SBFI and minBIC both underestimate high ranks.
-At high ranks, SBFI and minBIC perform best in terms of precision and sensitivity, meaning that the signatures our models do estimate are correct, even though some are missed, while SignatureAnalyzer yields some noisy or split signatures even if it learns rank correctly.
-For an Exponential / L1 prior (empty shapes), the sparsity penalty of SBFI may be too strong, as SBFI underestimates high ranks by a wide margin (up to 8 signatures under for $N=20$).
-However, minBIC still outperforms SignatureAnalyzer in terms of precision and sensitivity.
+For a Truncated Normal / L2 prior (filled shapes), all approaches estimate rank well, but with different error patterns: SignatureAnalyzer overestimates low and high ranks, while SBFI and minBIC underestimate high ranks.
+Under this prior and at high ranks, SBFI and minBIC achieve the highest precision and sensitivity, meaning that the signatures our models do estimate are correct, even though some are missed, while SignatureAnalyzer often yields noisy or split signatures, even if it learns the correct rank.
+For an Exponential / L1 prior (empty shapes), the sparsity penalty of SBFI may be too strong, as SBFI underestimates high ranks by a wide margin.
+In this setting, minBIC still shows higher precision and sensitivity than SignatureAnalyzer.
 With our Poisson+MH samplers, we are able to report posterior uncertainty at the cost of 12 additional minutes over SignatureAnalyzer (4 vs 16 minutes).
@@
 6.1 Data
-In this section, we analyze the Pan-Cancer Analysis of Whole Genomes (PCAWG) mutational counts data for whole genome sequencing (WGS) 96-alphabet single base substitution (SBS) mutation classification Consortium [2020].
-This is publicly available from the International Cancer Genome Consortium’s (ICGC) Accelerating Research in Genomic Oncology (ARGO) data portal [Zhang et al., 2019, access link].
+In this section, we analyze the Pan-Cancer Analysis of Whole Genomes (PCAWG) mutation count data using the 96-alphabet single base substitution (SBS) mutation classification [Consortium, 2020].
+This is publicly available from the International Cancer Genome Consortium’s (ICGC) Accelerating Research in Genomic Oncology (ARGO) data portal [Zhang et al., 2019].
 This includes 37 datasets, each for a different histology group.
@@
 For each histology group, we identify hypermutated samples by assuming the total mutation count per sample follows a Negative Binomial mixture model, with at least one component for non-hypermutated samples, and possibly one or more components for hypermutated samples (details in Appendix D.1).
-This removes 65 samples total (<<3% of data), with at most 13 samples removed per histology group in Skin-Melanoma.
+This removes 65 samples total (<<3% of data), with at most 13 samples removed per histology group in skin melanoma.
 6.2 Methods
-Because ground truth is unknown for real data applications, rigorous conclusions can only be made in the context of simulated data as in Section 5.
-However, we wish to illustrate our method and compare it to others in a real world setting.
-We compare our results to SignatureAnalyzer and show our approach results in similar biological interpretations.
-We use Poisson-Truncated Normal+MH SBFI as this specification is the closest to the Gamma priors used in most mutational signatures literature.
-We fit a separate sampler to each histology group, learning ranks $\in[1,20]$.
-For comparison, we also fit SignatureAnalyzer with a Poisson objective and L2 priors (equivalent to Truncated Normal priors).
-Estimated signatures are assigned to the COSMIC reference as described in Section 3.4.
+Because ground truth is unknown for real data applications, rigorous conclusions can only be made in the context of simulated data, as in Section 5.
+However, we wish to illustrate our method and compare it to others in a real-world setting.
+We compare our results to SignatureAnalyzer and show that our approach results in similar biological interpretations.
+We use the bayesNMF Poisson+MH approach with Truncated Normal priors and SBFI for rank $\in[1,20]$. For each histology group, we fit bayesNMF+MH, and compare to SignatureAnalyzer with L2 priors.
 Figure 4: Results of bayesNMF+SBFI (upwards triangle) and SignatureAnalyzer+ARD (downwards triangle) on PCAWG histology groups.
-Skin Melanoma is excluded for visual clarity because SignatureAnalyzer estimates a high rank of 76, (bayesNMF estimates a rank of 10).
-A) COSMIC reference signatures aligned to $\hat{P}$.
+Skin melanoma is excluded for visual clarity because SignatureAnalyzer estimates a high rank of 76 (bayesNMF estimates a rank of 10).
+A.
+COSMIC reference signatures aligned to $\hat{P}$.
 Reference signatures not found with either method are excluded.
-B) Latent ranks estimated by bayesNMF+SBFI and SignatureAnalyzer+ARD.
+B.
+Estimated latent ranks.
 Exact values in Appendix Table D.1.
@@
 [2020], including the presence of SBS1, SBS5, and SBS40 in nearly every cancer type.
-More detailed results including the median number of mutations attributed to each signature in each histology group by each method can be found in Appendix Figure D.2.
+More detailed results, including the median number of mutations attributed to each signature in each histology group by each method, can be found in Appendix Figure D.2.
 A clear difference between methods is the estimated rank—bayesNMF+SBFI typically results in a sparser solution, with SignatureAnalyzer+ARD frequently estimating twice the number of signatures (Figure 4B).
-Based on our simulation results, because bayesNMF+SBFI with Truncated Normal priors consistently achieved higher sensitivity and precision, we speculate that results here of SignatureAnalyzer with L2 priors may include spurious components (especially for Skin Melanoma, where SignatureAnalyzer estimates a rank of 76), making the resulting signatures less biologically interpretable.
+Based on our simulation results, because bayesNMF+SBFI with Truncated Normal priors consistently achieved higher sensitivity and precision, we speculate that the results here of SignatureAnalyzer with L2 priors may include spurious components (especially for skin melanoma, where SignatureAnalyzer estimates a rank of 76), making the resulting signatures less biologically interpretable.
 In most cases where bayesNMF recovered a signature not reported by SignatureAnalyzer (Figure 4A upwards triangles), the discrepancy can be partly explained by posterior uncertainty.
@@
 However, of the 34 signatures of this type, 28 (82%) had a SignatureAnalyzer-derived signature among the candidates receiving assignment votes, just not the majority required for posterior ensemble assignment.
-Therefore, the signature estimated by SignatureAnalyzer can be viewed as consistent with one plausible posterior mode inferred by bayesNMF.
-There are 6 signatures still unique to bayesNMF after considering this aspect, of which many match previous discoveries in the COSMIC database: SBS5 in CNS pilocytic astrocytoma and pancreatic adenocarcinoma, SBS1 in lymphocytic leukemia (CLL), and SBS95 in liver hepatocellular carcinoma (HCC).
+Therefore, many signatures estimated by SignatureAnalyzer can be viewed as consistent with one plausible posterior mode inferred by bayesNMF.
+Still, there are 6 signatures unique to bayesNMF after considering this aspect, of which many match previous discoveries in the COSMIC database: SBS5 in CNS pilocytic astrocytoma and pancreatic adenocarcinoma, SBS1 in lymphocytic leukemia (CLL), and SBS95 in liver hepatocellular carcinoma (HCC).
 Some are not in the COSMIC database, but have been previously reported in the literature: SBS36 in prostate adenocarcinoma [Jiang et al., 2020].
-However, there is no reference for SBS9 in esophageal adenocarcinoma, although it has been found in esophageal small cell carcinoma.
+There is no reference for SBS9 in esophageal adenocarcinoma, although it has been found in esophageal small cell carcinoma.
 7 Discussion
-In this paper, we introduced bayesNMF, a novel computationally efficient Gibbs sampler for Bayesian Poisson NMF and an automatic rank-learning approach.
-Through simulation studies and a large-scale data analysis, we demonstrated that bayesNMF achieves comparable accuracy and uncertainty quantification to standard Bayesian Poisson NMF at a fraction of the computational cost.
+We introduced bayesNMF, a MH-within-Gibbs sampler for Bayesian Poisson NMF with automatic rank-learning.
+The method is built on two primary innovations: an efficient MH-within-Gibbs sampling strategy to avoid Poisson augmentation, and a BIC-based sparsity prior that enables rank to be learned accurately within a single Bayesian formulation.
+Simulation studies and a large-scale data analysis demonstrate that bayesNMF achieves comparable accuracy and uncertainty quantification to standard Bayesian Poisson NMF at a fraction of the computational cost.
 All methods are available in an open-source R software package on GitHub at jennalandy/bayesNMF.
-The Metropolis-Hastings steps are the key to bayesNMF’s computational efficiency.
-In future work, these updates could easily be used in conjunction with other approaches, including compressive NMF [Zito and Miller, 2024] or multistudy NMF [Grabski et al., 2023b].
-This concept of model-paired proposals may also be generalizable to other hierarchical count models that rely on Poisson augmentation, for example Poisson regression with random effects.
+The MH steps with our high-overlap proposals are the key to bayesNMF’s computational efficiency.
+In future work, these updates could be used in conjunction with other approaches such as compressive NMF [Zito and Miller, 2024] or multistudy NMF [Grabski et al., 2025].
+Additionally, our model-paired proposals may be generalizable to other hierarchical count models that rely on Poisson augmentation, like Poisson regression with random effects.
 We recognize a potential limitation of our assumption that each estimated signature must be assigned to a unique reference signature.
-In some cases, precision reported in Figure 3 could be higher if multiple estimated signatures were allowed to map to the same reference.
-Similarly, some SignatureAnalyzer-estimated signatures in the PCAWG analysis that show cosine similarity to their assigned references (Appendix Figure D.2) may have achieved higher similarity if duplicate assignments were permitted.
-This is only a limitation of our downstream analysis and visualization tools, not of the models themselves.
+In some cases, the precision reported in Figure 3 could be higher if multiple estimated signatures were allowed to map to the same reference.
+Similarly, some SignatureAnalyzer-estimated signatures in the PCAWG analysis (Appendix Figure D.2) may have achieved higher similarity to their assigned references if duplicate assignments were permitted.
+However, this is only a limitation of our downstream analysis and visualization tools, not of the models themselves.
 In summary, bayesNMF offers a powerful toolbox for exploring complex mutational landscapes, especially with its ability to report posterior uncertainty while learning rank as part of the Bayesian model.
-The open source R software package encourages usability of these methods and reproducibility of our results.
-As cancer genomics databases expand, bayesNMF’s efficiency and flexibility make it well suited to drive further research.
+The open source R software package facilitates the usability of these methods and the reproducibility of our results.
+As cancer genomic databases expand, bayesNMF’s efficiency and flexibility make it well-suited to drive further research.
 8 Acknowledgments
@@
 Supplementary Material
-Appendices:
+Appendix:
 A) Model specifications (likelihoods, priors, hyperpriors) and Gibbs updates.
 B) Implementation details (convergence, warmup, inference, reference alignment, label switching diagnostic).
-C) Additional simulation study results, including sensitivity analyses for sparse data and changing rank range.
-D) Additional PCWAG data application results.
-(AppendixA.pdf, etc.)
+C) Additional simulation study results, including sensitivity analyses for sparse data and increasing rank range.
+D) Additional PCAWG data application results.
 R package:
-The R package bayesNMF is publicly available on GitHub at jennalandy/bayesNMF. Documented by README.md and vignettes.
+The R package bayesNMF is publicly available on GitHub at jennalandy/bayesNMF, documented by the README and vignettes.
 Code:
-All code needed to reproduce simulation studies and data analysis is publicly available on GitHub at jennalandy/bayesNMF_PAPER, documented by README.md.
+All code needed to reproduce simulation studies and data analysis is publicly available on GitHub at
+jennalandy/bayesNMF_PAPER, documented by README.md.
 Dataset:
-The PCAWG data is available through the ICGC ARGO data portal: access instructions link. Instructions are also included in the code file 1_pcawg_data.qmd.
+The PCAWG data are available through the ICGC ARGO data portal:
+access instructions at docs.icgc-argo.org/docs/data-access/. Instructions are also included in the code file 1_pcawg_data.qmd.
 Appendix A Appendix A: Model Specifications
@@
 Table A.3 summarizes the six models implemented in bayesNMF and links to the applicable subsection of this Appendix.
+We introduce all models using the simpler prior on the signature inclusion matrix $A$ corresponding to Bayesian Factor Inclusion (BFI).
+See the main text for details on the prior of $A$ under Sparse BFI (SBFI).
+If rank is fixed, $A$ is set as an identity matrix and Gibbs updates of $A$ are skipped.
 Throughout this appendix, let $\hat{M}$ be the reconstruction $\hat{P}\hat{A}\hat{E}$ such that $\hat{M}_{kg}=(\hat{P}\hat{E})_{kg}=\sum_{n}\hat{P}_{kn}\hat{E}_{ng}$.
@@
 The value $\bar{M}$ is the mean across all count values in the data matrix $M$, representing the general scale of the data.
-Likelihood Prior Sampler # Updates # Acceptance Ratios Section
+Likelihood Prior Sampler # Updates # Acc. Ratios Sec.
 Poisson TruncNormal MH-within-Gibbs $3(KN+NG)$ $KN+NG$ A.1
@@
 Typically $N<<K,G$.
-MH-within-Gibbs samplers are the new contribution of this work.
 A.0.1 Definitions Shared Across Models
@@
 No tempered samples are used for inference.
-A.1 Poisson - TruncNormal + MH
+A.1 Poisson - TruncNormal(TN) + MH
 A.1.1 Model Definition
 Likelihood: $M_{kg}\sim\text{Poisson}\left(\sum_{n}P_{kn}A_{nn}E_{ng}\right)$
-Priors: $P_{kn}\sim\text{TruncNormal}_{[0,\infty)}\left(\mu^{P}_{kn},(\sigma^{P}_{kn})^{2}\right)$
-$E_{ng}\sim\text{TruncNormal}_{[0,\infty)}\left(\mu^{E}_{ng},(\sigma^{E}_{ng})^{2}\right)$
+Priors: $P_{kn}\sim\text{TruncNorm}\left(\mu^{P}_{kn},(\sigma^{P}_{kn})^{2},0,\infty\right)$
+$E_{ng}\sim\text{TruncNorm}\left(\mu^{E}_{ng},(\sigma^{E}_{ng})^{2},0,\infty\right)$
 Hyperpriors: $\mu^{P}_{kn}\sim\text{Normal}\left(m^{P}_{kn}=0,(s^{2}_{kn})^{2}=\sqrt{\bar{M}/N}\right)$
@@
 The prior and hyperprior structure on $A$ is defined in Section A.0.1.
-From here on, let $\theta$ to refer to all prior and hyperprior parameters, and assume any use of “TruncNorm” uses the bounds $0,\infty)$.
+From here on, let $\theta$ to refer to all prior and hyperprior parameters, and assume the parameter order for Truncated Normal distributions $\text{TruncNorm}(\mu,\sigma^{2},a,b)$ where $a$ and $b$ are truncation bounds.
 This model has $KN+NG$ parameters and $2(KN+NG)$ prior parameters to update on each iteration. MH-within-Gibbs steps are used for $P$ and $E$, meaning $KN+NG$ acceptance ratios must be computed.
@@
 Targets
-$f(P^{*}_{kn}|P,E,M,\theta)$ $\propto p_{\text{TruncNorm}}\left(P^{*}_{kn}|\mu^{P}_{kn},(\sigma^{P}_{kn})^{2}\right)\prod_{g}p_{\text{Poisson}}\left(M_{kg}|\lambda=(P^{*}E)_{kg}\right)$
-$f(E^{*}_{ng}|P,E,M,\theta)$ $\propto p_{\text{TruncNorm}}\left(E^{*}_{ng}|\mu^{E}_{ng},(\sigma^{E}_{ng})^{2}\right)\prod_{k}p_{\text{Poisson}}\left(M_{kg}|\lambda=(PE^{*})_{kg}\right)$
+$f(P^{*}_{kn}|P,E,M,\theta)$ $\propto p_{\text{TruncNorm}}\left(P^{*}_{kn}|\mu^{P}_{kn},(\sigma^{P}_{kn})^{2},0,\infty\right)\prod_{g}p_{\text{Poisson}}\left(M_{kg}|\lambda=(P^{*}E)_{kg}\right)$
+$f(E^{*}_{ng}|P,E,M,\theta)$ $\propto p_{\text{TruncNorm}}\left(E^{*}_{ng}|\mu^{E}_{ng},(\sigma^{E}_{ng})^{2},0,\infty\right)\prod_{k}p_{\text{Poisson}}\left(M_{kg}|\lambda=(PE^{*})_{kg}\right)$
 Proposals
-$g(P^{*}_{kn}|P,E,M,\theta)$ $\propto p_{\text{TruncNorm}}(P^{*}_{kn}|\mu^{P}_{kn},(\sigma^{P}_{kn})^{2})\prod_{g}p_{\text{Normal}}\left(M_{kg}|\mu=(P^{*}E)_{kg},\sigma^{2}=(PE)_{kg}\right)$
-$\sim\text{TruncNorm}_{[0,\infty)}\left(\mu=m^{P}_{kn}(M,P,A,E,\theta),\sigma^{2}=s^{P}_{kn}(M,P,E,\theta)\right)$
+$g(P^{*}_{kn}|P,E,M,\theta)$ $\propto p_{\text{TruncNorm}}(P^{*}_{kn}|\mu^{P}_{kn},(\sigma^{P}_{kn})^{2},0,\infty)\prod_{g}p_{\text{Normal}}\left(M_{kg}|\mu=(P^{*}E)_{kg},\sigma^{2}=(PE)_{kg}\right)$
+$P^{*}_{kn}|P,E,M,\theta$ $\sim\text{TruncNorm}\left(\mu=m^{P}_{kn}(M,P,A,E,\theta),\sigma^{2}=s^{P}_{kn}(M,P,E,\theta),0,\infty\right)$
 $m^{P}_{kn}(M,P,A,E,\theta)=\frac{\mu^{P}_{kn}/(\sigma^{P}_{kn})^{2}+\sum_{g=1}^{G}A_{nn}E_{ng}(M_{kg}-\hat{M}_{kg/n})/(PE)_{kg}}{1/(\sigma^{P}_{kn})^{2}+\sum_{g=1}^{G}A_{nn}E_{ng}^{2}/(PE)_{kg}}$
 $s^{P}_{kn}(M,P,A,E,\theta)=\frac{1}{1/(\sigma^{P}_{kn})^{2}+\sum_{g=1}^{G}A_{nn}E_{ng}^{2}/(PE)_{kg}}$
-$g(E^{*}_{ng}|P,E,M,\theta)$ $\propto p_{\text{TruncNorm}}(E^{*}_{ng}|\mu^{E}_{ng},(\sigma^{E}_{ng})^{2})\prod_{k}p_{\text{Normal}}\left(M_{kg}|\mu=(PE^{*})_{kg},\sigma^{2}=(PE)_{kg}\right)$
-$\sim\text{TruncNorm}_{[0,\infty)}\left(\mu=m^{E}_{ng}(M,P,A,E,\theta),\sigma^{2}=s^{E}_{ng}(M,P,A,E,\theta)\right)$
+$g(E^{*}_{ng}|P,E,M,\theta)$ $\propto p_{\text{TruncNorm}}(E^{*}_{ng}|\mu^{E}_{ng},(\sigma^{E}_{ng})^{2},0,\infty)\prod_{k}p_{\text{Normal}}\left(M_{kg}|\mu=(PE^{*})_{kg},\sigma^{2}=(PE)_{kg}\right)$
+$E^{*}_{ng}|P,E,M,\theta$ $\sim\text{TruncNorm}\left(\mu=m^{E}_{ng}(M,P,A,E,\theta),\sigma^{2}=s^{E}_{ng}(M,P,A,E,\theta),0,\infty\right)$
 $m^{E}_{ng}(M,P,A,E,\theta)=\frac{\mu^{E}_{ng}/(\sigma^{E}_{ng})^{2}+\sum_{k=1}^{K}P_{kn}A_{nn}(M_{kg}-\hat{M}_{kg/n})/(PE)_{kg}}{1/(\sigma^{E}_{ng})^{2}+\sum_{k=1}^{K}P_{kn}^{2}A_{nn}/(PE)_{kg}}$
@@
 $g(P^{*}_{kn}|P,E,M,\theta)$ $\propto p_{\text{Exp}}(P^{*}_{kn}|\lambda^{P}_{kn})\prod_{g}p_{\text{Normal}}\left(M_{kg}|\mu=(P^{*}E)_{kg},\sigma^{2}=(PE)_{kg}\right)$
-$\sim\text{TruncNorm}_{[0,\infty)}\left(\mu=m^{P}_{kn}(M,P,A,E,\theta),\sigma^{2}=s^{P}_{kn}(M,P,E,\theta)\right)$
+$P^{*}_{kn}|P,E,M,\theta$ $\sim\text{TruncNorm}\left(\mu=m^{P}_{kn}(M,P,A,E,\theta),\sigma^{2}=s^{P}_{kn}(M,P,E,\theta),0,\infty\right)$
 $m^{P}_{kn}(M,P,A,E,\theta)=\frac{\sum_{g=1}^{G}A_{nn}E_{ng}(M_{kg}-\hat{M}_{kg/n})/(PE)_{kg}-\lambda^{P}_{kn}}{\sum_{g=1}^{G}A_{nn}E_{ng}^{2}/(PE)_{kg}}$
@@
 $g(E^{*}_{ng}|P,E,M,\theta)$ $\propto p_{\text{Exp}}(E^{*}_{ng}|\lambda^{E}_{ng})\prod_{k}p_{\text{Normal}}\left(M_{kg}|\mu=(PE^{*})_{kg},\sigma^{2}=(PE)_{kg}\right)$
-$\sim\text{TruncNorm}_{[0,\infty)}\left(\mu=m^{E}_{ng}(M,P,A,E,\theta),\sigma^{2}=s^{E}_{ng}(M,P,A,E,\theta)\right)$
+$E^{*}_{ng}|P,E,M,\theta$ $\sim\text{TruncNorm}\left(\mu=m^{E}_{ng}(M,P,A,E,\theta),\sigma^{2}=s^{E}_{ng}(M,P,A,E,\theta),0,\infty\right)$
 $m^{E}_{ng}(M,P,A,E,\theta)=\frac{\sum_{k=1}^{K}P_{kn}A_{nn}(M_{kg}-\hat{M}_{kg/n})/(PE)_{kg}-\lambda^{E}_{ng}}{\sum_{k=1}^{K}P_{kn}^{2}A_{nn}/(PE)_{kg}}$
@@
 The update for $A_{nn}$ is defined in Section A.0.2.
-A.5 Normal - TruncNormal
+A.5 Normal - TruncNormal (TN)
 A.5.1 Model Definition
 Likelihood: $M_{kg}\sim\text{Normal}\left(\sum_{n}P_{kn}A_{nn}E_{ng},\sigma^{2}_{g}\right)$
-Priors: $P_{kn}\sim\text{TruncNormal}_{[0,\infty)}\left(\mu^{P}_{kn},(\sigma^{P}_{kn})^{2}\right)$
-$E_{ng}\sim\text{TruncNormal}_{[0,\infty)}\left(\mu^{E}_{ng},(\sigma^{E}_{ng})^{2}\right)$
+Priors: $P_{kn}\sim\text{TruncNorm}\left(\mu^{P}_{kn},(\sigma^{P}_{kn})^{2},0,\infty\right)$
+$E_{ng}\sim\text{TruncNorm}\left(\mu^{E}_{ng},(\sigma^{E}_{ng})^{2},0,\infty\right)$
 $\sigma^{2}_{g}\sim\text{InverseGamma}(\alpha_{g}=3,\beta_{g}=3)$
@@
 The prior and hyperprior structure on $A$ is defined in Section A.0.1.
-From here on, let $\theta$ to refer to all prior and hyperprior parameters, and assume any use of “TruncNorm” uses the bounds $0,\infty)$.
+From here on, let $\theta$ to refer to all prior and hyperprior parameters, and assume the parameter order for Truncated Normal distributions $\text{TruncNorm}(\mu,\sigma^{2},a,b)$ where $a$ and $b$ are truncation bounds.
 This model has $KN+NG+G$ parameters and $2(KN+NG)$ prior parameters to update on each iteration.
@@
 A.5.3 Parameter Gibbs Updates
-$P_{kn}|$... $\sim\text{TruncNorm}_{[0,\infty)}\left(\mu=m^{P}_{kn}(M,A,E,\theta),\sigma^{2}=s^{P}_{kn}(M,A,E,\theta)\right)$
+$P_{kn}|$... $\sim\text{TruncNorm}\left(\mu=m^{P}_{kn}(M,A,E,\theta),\sigma^{2}=s^{P}_{kn}(M,A,E,\theta),0,\infty\right)$
 $m^{P}_{kn}(M,A,E,\theta)=\frac{\mu^{P}_{kn}/(\sigma^{P}_{kn})^{2}+\sum_{g=1}^{G}A_{nn}E_{ng}(M_{kg}-\hat{M}_{kg/n})/\sigma^{2}_{g}}{1/(\sigma^{P}_{kn})^{2}+\sum_{g=1}^{G}A_{nn}E_{ng}^{2}/\sigma^{2}_{g}}$
 $s^{P}_{kn}(M,A,E,\theta)=\frac{1}{1/(\sigma^{P}_{kn})^{2}+\sum_{g=1}^{G}A_{nn}E_{ng}^{2}/\sigma^{2}_{g}}$
-$E_{ng}|$... $\sim\text{TruncNorm}_{[0,\infty)}\left(\mu=m^{E}_{ng}(M,P,A,\theta),\sigma^{2}=s^{E}_{ng}(M,P,A,\theta)\right)$
+$E_{ng}|$... $\sim\text{TruncNorm}\left(\mu=m^{E}_{ng}(M,P,A,\theta),\sigma^{2}=s^{E}_{ng}(M,P,A,\theta),0,\infty\right)$
 $m^{E}_{ng}(M,P,A,\theta)=\frac{\mu^{E}_{ng}/(\sigma^{E}_{ng})^{2}+\sum_{k=1}^{K}P_{kn}A_{nn}(M_{kg}-\hat{M}_{kg/n})/\sigma^{2}_{g}}{1/(\sigma^{E}_{ng})^{2}+\sum_{k=1}^{K}P_{kn}^{2}A_{nn}/\sigma^{2}_{g}}$
@@
 A.6.3 Parameter Gibbs Updates
-$P_{kn}$... $\sim\text{TruncNorm}_{[0,\infty)}\left(\mu=m^{P}_{kn}(M,A,E,\theta),\sigma^{2}=s^{P}_{kn}(M,A,E,\theta)\right)$
+$P_{kn}$... $\sim\text{TruncNorm}\left(\mu=m^{P}_{kn}(M,A,E,\theta),\sigma^{2}=s^{P}_{kn}(M,A,E,\theta),0,\infty\right)$
 $m^{P}_{kn}(M,A,E,\theta)=\frac{\sum_{g=1}^{G}A_{nn}E_{ng}(M_{kg}-\hat{M}_{kg/n})/\sigma^{2}_{g}-\lambda^{P}_{kn}}{\sum_{g=1}^{G}A_{nn}E_{ng}^{2}/\sigma^{2}_{g}}$
 $s^{P}_{kn}(M,A,E,\theta)=\frac{1}{\sum_{g=1}^{G}A_{nn}E_{ng}^{2}/\sigma^{2}_{g}}$
-$E_{ng}|$... $\sim\text{TruncNorm}_{[0,\infty)}\left(\mu=m^{E}_{ng}(M,P,A,\theta),\sigma^{2}=s^{E}_{ng}(M,P,A,\theta)\right)$
+$E_{ng}|$... $\sim\text{TruncNorm}\left(\mu=m^{E}_{ng}(M,P,A,\theta),\sigma^{2}=s^{E}_{ng}(M,P,A,\theta),0,\infty\right)$
 $m^{E}_{ng}(M,P,A,\theta)=\frac{\sum_{k=1}^{K}P_{kn}A_{nn}(M_{kg}-\hat{M}_{kg/n})/\sigma^{2}_{g}-\lambda^{E}_{ng}}{\sum_{k=1}^{K}P_{kn}^{2}A_{nn}/\sigma^{2}_{g}}$
@@
 We use the following approach to align estimated signatures to ground truth in our simulation studies as well.
-We follow existing mutational signatures works [Rosales et al., 2016, Gori and Baez-Ortega, 2018, Alexandrov et al., 2020, Grabski et al., 2023b] and focus on the scale-independent measure of cosine similarities between signatures.
+We follow existing mutational signatures works [Rosales et al., 2016, Gori and Baez-Ortega, 2018, Alexandrov et al., 2020, Grabski et al., 2025] and focus on the scale-independent measure of cosine similarities between signatures.
 To align estimated factors to the reference factors, we first create a similarity matrix $S$, where each element is the cosine similarity between a column of $P$ and a column of $\hat{P}$, $S_{ij}=cos(P_{i},\hat{P}_{j})$.
@@
 D.2 Additional Results
-For PCAWG results using bayesNMF+SBFI and SignatureAnalyzer+ARD, Figure D.2 reports median number of mutations attributed to each signature and the cosine similarity of each estimated signature to its assigned reference signature (posterior average for bayesNMF).
-Table D.5 reports estimated rank for PCAWG results using bayesNMF+SBFI and SignatureAnalyzer+ARD.
-Figure D.3 investigates signatures assigned by both bayesNMF and SignatureAnalyzer versus signatures assigned by bayesNMF only.
+  • Figure D.2: For PCAWG results using bayesNMF+SBFI and SignatureAnalyzer+ARD, this figure reports median number of mutations attributed to each signature and the cosine similarity of each estimated signature to its assigned reference signature (posterior average for bayesNMF).
+  • Table D.5: Reports estimated rank for PCAWG results using bayesNMF+SBFI and SignatureAnalyzer+ARD.
+  • Figure D.3: Investigates signatures assigned by both bayesNMF and SignatureAnalyzer versus signatures assigned by bayesNMF only.
 The signatures assigned by bayesNMF only have lower cosine similarity to the assigned reference and a lower proportion of assignment votes, indicating uncertainty in assignment.